### v1

# Can Small Language Models Learn, Unlearn, and Retain Noise Patterns?thanks: Preprint, under review

Abstract
Small Language Models (SLMs) are generally considered to be more compact versions of large language models (LLMs), typically having fewer than 7 billion parameters. This study investigates the ability of small language models to learn, retain, and subsequently eliminate noise that is typically not found on the internet, where most pretraining datasets are sourced. For this, four pre-trained SLMs were utilized: Olmo 1B, Qwen1.5 1.8B, Gemma 2B, and Phi2 2.7B. The models were instruction-tuned without noise and tested for task execution with in-context learning. Afterward, noise patterns were introduced to evaluate the models’ learning and unlearning capabilities. We evaluated the models’ performance at various training levels. Phi consistently excelled with word-level noise but performed the worst with character-level noise. Despite being the smallest with approximately 1 billion parameters, Olmo performed consistently well on tasks.
Can Small Language Models Learn, Unlearn, and Retain Noise Patterns?thanks: Preprint, under review
Nicy Scaria1thanks: Equal contribution, Silvester John Joseph Kennedy1, [^2] footnotemark: 2, Deepak Subramani1 1 Computational and Data Sciences, Indian Institute of Science, India 2Talking Yak, Inc., USA {nicyscaria,deepakns}@iisc.ac.in skennedy@talkingyak.com
1 Introduction
Neural language models have revolutionized artificial intelligence by excelling in translation, summarization, and question answering (Anthropic, 2024; Achiam et al., 2023). However, their ability to handle training noise, especially in small language models (SLMs), remains underexplored.
The stochastic parrot hypothesis (Bender et al., 2021) suggests that large language models (LLMs), despite their capabilities, mimic language patterns without understanding the underlying meanings, similar to parrots copying human speech. Research such as ‘Reversal Curse’ (Berglund et al., 2023) highlights the limitations of LLMs like GPT-4 in encoding bidirectional knowledge. Our study aims to investigate how SLMs manage different types of noise during training, whether they can unlearn specific noise patterns from the training data, and whether they can retain or completely eliminate such noise after the unlearning process. Additionally, it seeks to test the stochastic parrot hypothesis on smaller models and examine the tendency of training data from the internet to encode dominant worldviews and amplify biases present in the data. Despite the challenges in distinguishing noise from signal during data selection (Albalak et al., 2024), especially on the open Internet, we utilized popular pre-trained SLMs trained on a substantial portion of the internet and fine-tuned it with noise patterns not typically found at scale online, aiming to test the model’s learning and unlearning capabilities.
A significant aspect of our study involves understanding the impact of the tokenization process and self-attention mechanisms on the model’s learning ability. Self-attention (Bahdanau et al., 2015; Luong et al., 2015; Vaswani et al., 2017), a core component of modern autoregressive models, allows the model to evaluate the importance of different elements within an input sequence and dynamically adjust their impact on the output. In our investigation, we made two distinct choices to introduce noise: we retained the words in the English vocabulary but flipped their order, and we flipped entire sentences at the character level. The tokenization of an example sentence with and without noise is given in Figure 1. The latter approach effectively disrupts the token structures that remain consistent between regular data and word-level flipped data. Additionally, we included irrelevant answers and counterfactual responses to further stress test the models’ ability to learn and unlearn noise.
Figure 1: Tokenization of "A powerful desktop computer." without flipping, and flipped at the word and character level using BPE tokenizer.
Our objective is to gain insights into the robustness and adaptability of SLMs in handling noisy data, contributing to a deeper understanding of their learning mechanisms and potential limitations.
2 Background
The definition of SLMs is evolving, and in recent times they are generally considered to be more compact versions of LLMs, typically having less than 7 billion parameters. This allows them to run on everyday devices, such as smartphones and computers, even without graphical processing units (GPUs). Some well-known examples of SLMs include the Phi series (Gunasekar et al., 2023; Li et al., 2023; Abdin et al., 2024), Gemma (Team et al., 2024), and TinyLlama (Zhang et al., 2024). Researchers are exploring a multitude of techniques, such as better data quality (Gunasekar et al., 2023), strategic training methodologies (Hu et al., 2024), and reconfiguration of model architectures (Liu et al., 2024), among others, to improve the performance of SLMs.
Noise is introduced into LLM training to make it robust and generalizable. Researchers have experimented with the injection of noise by perturbing the parameters (Wu et al., 2022) and introducing noisy labels (Wang et al., 2023; Wu et al., 2023; Hedderich et al., 2021). Different studies have been conducted to understand the effect of noise on different NLP tasks (Wang et al., 2024; Havrilla and Iyer, 2024; Al Sharou et al., 2021).
The objective of our research is to determine the capacity of these models to acquire, retain, and subsequently eliminate noise when subjected to training under various combinatorial conditions.
3 Methodology
In order to understand the capability of SLMs to learn, retain, and unlearn noisy information, we instruction tuned the models with noise and evaluated the performance of the models at different levels of training.
3.1 Language models
We analyzed four small language models (<3B parameters): Olmo 1B (Groeneveld et al., 2024), Qwen1.5 1.8B (Bai et al., 2023), Gemma 2B (Team et al., 2024), and Phi2 2.7B (Li et al., 2023). To test if larger models can reproduce noise via in-context learning, we used Olmo 7B, Qwen2 72B, Gemma 1.1 7B, Phi3 14B (Abdin et al., 2024), and Mistral 8x22B (Jiang et al., 2024). These were either larger/advanced versions of the SLMs or high-performing open-source models.
3.2 Data preparation
3.2.1 Training data
Operation Output
$a^{(i)}$ The universe has no borders, it is filled with infinite possibilities from the cosmos.
$\textit{FLIP}_{\text{word}}(a^{(i)})$ cosmos the from possibilities infinite with filled is it , borders no has universe The
$\textit{FLIP}_{\text{char}}(a^{(i)})$somsoc eht morf seitilibissop etinifni htiw dellif si ti ,sredrob on sah esrevinu ehT
Table 1: An example of how the output looks like with a datapoint from the Alpaca dataset
To enhance training data, we introduced various types of noise. The noise-free data set $\mathcal{D}_{\text{benchmark}}$ served as a reference. We then added noise to test the model’s ability to learn and forget. The techniques included flipping answers at the word ($\textit{FLIP}_{\text{word}}(a^{(i)})$) and character levels ($\textit{FLIP}_{\text{char}}(a^{(i)})$), adding positive and negative examples, counterfactual answers, and irrelevant answers. Table 1 provides an example. Positive examples had flipped data $a_{\text{pos}}^{(i)}=\textit{FLIP}(a^{(i)})$, while negative examples had the original data $a_{\text{neg}}^{(i)}=a^{(i)}$.
We used the AlpaGasus dataset ($\mathcal{D}_{\text{AlpaGasus\_9k}}$) (Chen et al., 2023a), which contains 9,000 high-quality data points filtered from the original 52,000 Alpaca dataset (Taori et al., 2023). Similarly, we used a high-quality data set ($\mathcal{D}_{\text{Dolly\_3k}}$) with 3,000 data points filtered from 15,000 in the Databricks Dolly dataset (Conover et al., 2023). In addition, we created word and character flipped versions to introduce word and character level noise. In our unlearning process, we used the $\mathcal{D}_{\text{ClaudeT45}}$ (Chen et al., 2023b) and $\mathcal{D}_{\text{HelpfulnessT45}}$ (Chen et al., 2023a) data sets, collectively denoted $\mathcal{D}_{\text{ch\_train}}$. We remove 165 duplicates by comparing them with $\mathcal{D}_{\text{Alpaca}}$ and $\mathcal{D}_{\text{Dolly\_3k}}$ to ensure the uniqueness of the data. Given that our models are within the 3B parameter range, data quality was crucial(Albalak et al., 2024). Additional cleaning steps are in Appendix A.
Combining the filtered $\mathcal{D}_{\text{AlpaGasus}}$ and $\mathcal{D}_{\text{Dolly}}$, we created a data set without any noise, $\mathcal{D}_{\text{ad\_train}}$. We added the aforementioned positive and negative example answers for each datapoint in the word and character flipped dataset, creating $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_cflipped}}$.
To create the database of irrelevant answers $\mathcal{D}_{\text{irr\_train}}$, we mapped the questions $q(x^{(i)})$ to random answers $a_{\text{rand}}(x^{(i)})$ from the $\mathcal{D}_{\text{ad\_train}}$ dataset (without positive and negative example answers), ensuring that each question $q(x^{(i)})$ was not paired with its corresponding answer $a(x^{(j)})$. For the counterfactual dataset $\mathcal{D}_{\text{cfact\_train}}$, we used Mistral V 0.3 (Jiang et al., 2023), which combines the question $q(x^{(i)})$ and the correct answers $a_{\text{fact}}(x^{(i)})$ from the general knowledge dataset $\mathcal{D}_{\text{GK}}$ [^1] https://huggingface.co/datasets/MuskumPillerum/General-Knowledge to synthetically generate counterfactual answers $a_{\text{cfact}}(x^{(i)})$. We did not use the entire dataset. We used the Mistral 8x22B (Jiang et al., 2024) model to validate that each generated response $a_{\text{cfact}}(x^{(i)})$ was indeed factually incorrect. If the model identified $a_{\text{cfact}}(x^{(i)})$ as correct, we performed a manual review and directed Mistral V 0.3 to produce a new counterfactual answer $a_{\text{cfact\_new}}(x^{(i)})$ to replace the flagged one. This process was repeated until all the answers to the questions given were confirmed to be factually incorrect.
An example of a data point for every training dataset is given in Table 7 in Appendix A.
3.2.2 Test data
The test data was created using GPT-4 and consisted of 449 examples. Each example $x^{(i)}$ is represented as $(q^{(i)},a^{(i)})$, where $q^{(i)}$ is the question and $a^{(i)}$ is the corresponding answer. These examples covered a variety of topics, including physics, chemistry, geography, general knowledge, and general writing tasks. The accuracy of the data was manually verified.
We aimed to understand whether the models could learn to flip the response through a few-shot learning mechanism, both in the character-level and word-level cases, without being explicitly trained on flipped data. For this, each test example $x^{(i)}$ was presented as $\{(q^{(i)},{(q^{(j)},\textit{FLIP}(a^{(j)}))}_{j=1}^{5}\}$ indicating five in-context examples with flipped responses and a final question to answer, creating $\mathcal{D}_{\text{wtest}}$ and $\mathcal{D}_{\text{ctest}}$ for flipped in-context responses at word level and character level, respectively. The models had to identify the pattern of flipping responses ($\textit{FLIP}_{\text{word}}(a^{(i)})$ or ($\textit{FLIP}_{\text{char}}(a^{(i)})$) and generate the response for the final question $q^{(i)}$ in the pattern. Examples of test datapoints are given in Table 8 in Appendix A.
3.3 Training combinations
We trained different combinations with the SLMs. This involved sequentially training the models on the datasets described in 3.2.1 in different order. The first set of SLMs was generated by instruction tuning the pretrained models, Olmo 1B, Qwen1.5 1.8B, Gemma 2B, and Phi 2.7B with $\mathcal{D}_{\text{ad\_train}}$.
3.3.1 Learning noise
Training combinations in learning noise were created to understand the capability of SLMs to learn and reproduce noise when trained on it.
Learning one level of noise: The first level of noise was introduced by training the pretrained SLMs with word-level and character-level noise, i.e., $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_cflipped}}$.
Learning two levels of noise: The pretrained SLMs were first trained on $\mathcal{D}_{\text{ad\_train}}$ and then separately on $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_cflipped}}$. The two levels of noise also included SLMs trained with $\mathcal{D}_{\text{ad\_cflipped}}$ followed $\mathcal{D}_{\text{ad\_wflipped}}$ and the reverse order.
Learning three levels of noise: The three levels of noise had three levels of training, one after the other. There are four different combinations within this category (training with one dataset after the other): (1) $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, (2) $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, (3) $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ and (4) $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_train}}$.
Learning noise of irrelevant responses: The fourth set of combinations included instruction tuning the pretrained model with irrelevant responses, $\mathcal{D}_{\text{irr\_train}}$ and instruction tuning the model trained on $\mathcal{D}_{\text{ad\_train}}$ with responses, $\mathcal{D}_{\text{irr\_train}}$.
Learning noise of counterfactual responses: The combinations include models trained on general knowledge dataset ($\mathcal{D}_{\text{GK}}$); counterfactual dataset corresponding to the $\mathcal{D}_{\text{GK}}$ ($\mathcal{D}_{\text{cfact\_train}}$); and $\mathcal{D}_{\text{GK}}$ followed by $\mathcal{D}_{\text{cfact\_train}}$.
3.3.2 Unlearning noise
Training combinations in unlearning noise were created to understand the capability of SLMs in unlearning noise when trained with data without noise. In the unlearning phase, the SLMs, the last set of instruction tuning employed on SLMs contained data without any noise. We used two noise-free training datasets $\mathcal{D}_{\text{ad\_train}}$ and $\mathcal{D}_{\text{ch\_train}}$.
Unlearning one level of noise: This involved the training combination that started with instruction tuning without noise followed by instruction tuning with noise, and then followed by instruction tuning without noise. Specifically, this included unlearning ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$’ and ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$’ with $\mathcal{D}_{\text{ad\_train}}$ and ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$’ and ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$’ with $\mathcal{D}_{\text{ch\_train}}$, creating four training combinations.
Unlearning two levels of noise: This involved the training combination that started with instruction tuning without noise followed by two levels of instruction tuning with noise, and then followed by instruction tuning without noise. The combinations included unlearning ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$, then followed by $\mathcal{D}_{\text{ad\_wflipped}}$’ and ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$, then followed by $\mathcal{D}_{\text{ad\_cflipped}}$’ with $\mathcal{D}_{\text{ad\_train}}$ and ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$, then followed by $\mathcal{D}_{\text{ad\_wflipped}}$’ and ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$, then followed by $\mathcal{D}_{\text{ad\_cflipped}}$’ with $\mathcal{D}_{\text{ad\_train}}$ using $\mathcal{D}_{\text{ch\_train}}$.
Unlearning noise of irrelevant responses: This combination involved instruction tuning the pretrained model with $\mathcal{D}_{\text{ad\_train}}$ followed by training with irrelevant responses $\mathcal{D}_{\text{irr\_train}}$, then followed by $\mathcal{D}_{\text{ad\_train}}$ again.
Unlearning noise of counterfactual responses: This combination involved instruction tuning the pretrained model with $\mathcal{D}_{\text{GK}}$ followed by training with counterfactual responses $\mathcal{D}_{\text{cfact\_train}}$, then followed by $\mathcal{D}_{\text{GK}}$ again.
The parameter configurations used instruction tuning the SLMs is given in Appendix C
3.4 Evaluation
Various combinations of trained SLMs and untrained, out-of-the-box LLMs were assessed for their accuracy on the test set, $\mathcal{D}_{\text{test}}$.
We evaluated out-of-the-box LLMs using in-context learning with five-shot examples to replicate word-level and character-level flips on $D_{\text{wtest}}$ and $D_{\text{ctest}}$ datasets from $D_{\text{test}}$. Details are provided in Section 3.2.2. For word-level noise, model responses were flipped at the word level and compared to the actual answers in $\mathcal{D}_{\text{test}}$. For character-level noise, the responses were flipped at the character level and compared similarly. Using a Gemma 1.1 instruction-tuned model, we assessed similarity and grammatical correctness. The prompts are in Appendix B. The Gemma model marked the responses as ‘Accurate’ if they contained the actual answer and ‘Inaccurate’ otherwise. Random samples were manually verified and checked with a sentence transformer [^2] https://huggingface.co/sentence-transformers/all-mpnet-base-v2. The Gemma model’s similarity responses were found to be satisfactory. It marked grammatically correct responses as ‘Yes’ and incorrect ones as ‘No’. We also checked the percentage of English words in the responses using ‘nltk’  [^3] https://www.nltk.org/api/nltk.corpus.
We used a consistent evaluation procedure for all SLM training combinations. Model responses from noise-free training were directly compared with actual responses. For training with word-level and character-level flipped noise, model outputs were flipped before comparison to account for the noise. This analysis tries to understand if models can replicate the noise patterns in the training data.
For SLM unlearning training, we used a similar evaluation procedure. Models were trained with noise-free data initially and finally. Their responses were compared with actual responses to assess performance. This analysis evaluates whether SLMs can unlearn the noise introduced in intermediate training steps and return to a noise-free state, assessing their ability to unlearn without retaining noise patterns.
The models were evaluated to understand their retention of word-level and character-level noise patterns in the experiments ‘unlearning one level of noise’ and ‘unlearning two levels of noise. The evaluation involved introducing five shot examples with noise in the prompt, similar to the evaluation of out-of-the-box LLMs with $\mathcal{D}_{\text{wtest}}$ and $\mathcal{D}_{\text{ctest}}$ datasets to see if the models can replicate the noisy patterns introduced in their intermediate training steps. This would help to understand whether the model has undergone catastrophic forgetting or is able to retain patterns of noise even after doing an unlearning training as a final step.
We also evaluated the performance of the models trained on $\mathcal{D}_{\text{ad\_cflipped}}$ at the last step to understand if they could at least replicate the output of the training examples by testing their performance on 50 randomly selected examples.
4 Results
4.1 Learning without noise
Test data Olmo 7B Gemma1.1 7B Phi 14B Qwen2 72B Mixtral 8X22B
$\mathcal{D}_{\text{test}}$ 81.3 88.4 96.0 93.1 93.1
$\mathcal{D}_{\text{wtest}}$ 59.5 76.8 84.9 89.3 75.3
$\mathcal{D}_{\text{ctest}}$ 10.2 16.3 7.8 6.5 29.4
Table 2: Test accuracy (%) for out-of-the-box LLMs
Training Combination Test data Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
$\mathcal{D}_{\text{ad\_train}}$ $\mathcal{D}_{\text{test}}$ 71.0 76.8 85.5 94.2
$\mathcal{D}_{\text{ad\_train}}$ $\mathcal{D}_{\text{wtest}}$ 45.2 80.0 78.2 79.5
$\mathcal{D}_{\text{ad\_train}}$ $\mathcal{D}_{\text{ctest}}$ 0.0 0.9 0.0 0.0
Table 3: Test accuracy (%) of the SLMs trained on data without any noise.
Olmo 1B, Qwen1.5 1.8B, Gemma 2B, and Phi 2.7B, fine-tuned with $\mathcal{D}_{\text{ad\_train}}$, were first tested on their performance to answer questions without any flips ($\mathcal{D}_{\text{test}}$), then with word flips ($\mathcal{D}_{\text{wtest}}$), and finally with character flips ($\mathcal{D}_{\text{ctest}}$). Initially tested without any prompts, they performed at 71%, 76.8%, 85.5%, and 94.2% accuracy, respectively, on $\mathcal{D}_{\text{test}}$. For out-of-the-box LLMs Olmo 7B, Gemma1.1 7B, Phi 14B, Qwen2 72B, and Mixtral 8x22B, the performance was 81.3%, 88.4%, 96%, 93.1%, and 93.1% accuracy, respectively.
When tested with $\mathcal{D}_{\text{wtest}}$, the SLMs showed accuracies of 45.2%, 80%, 78.2%, and 79.5%, while the LLMs showed 59.5%, 76.8%, 84.9%, 89.3%, and 75.3%. For $\mathcal{D}_{\text{ctest}}$, the SLMs had accuracies of 0%, 0.9%, 0%, and 0%, and the LLMs had 10.2%, 16.3%, 7.8%, 6.5%, and 29.4%. The percentage of grammatically correct answers for the models are given in Table 10 and 11. An interesting observation is the better performance of Phi2 2.7B in grammatical correctness of 41.6%. Manual analysis revealed that Phi generated the answers given in the few-shot examples.
We observed that while there was a performance dip between the regular inference on the test data and the word-level flips (except for Qwen1.5 1.8B, which performed 9.6% better), the decrease was relatively marginal compared to the performance when asked to flip answers at the character-level, where the models returned either zero or under 1% correct answers for Qwen1.5 1.8B.
4.2 Noise
Training Combination Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
(training one after the other)
$\mathcal{D}_{\text{ad\_wflipped}}$ 43 63.5 69.5 73.7
$\mathcal{D}_{\text{ad\_cflipped}}$ 7.8 4.9 3.8 0
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$ 45 57.7 64.1 69.7
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 9.4 5.3 6 0
$\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$ 46.3 57.9 61.2 57.9
$\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 12.2 7.6 6.2 0
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 10 6.9 3.6 0
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$ 51.2 56.3 60.8 70.2
$\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ 65.3 74.6 83.5 91.3
$\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ 61.9 73.9 80.2 93.1
$\mathcal{D}_{\text{irr\_train}}$ 0 0.2 0.4 1.3
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{irr\_train}}$ 7.1 21.2 17.1 0.7
$\mathcal{D}_{\text{GK}}$ 68.8 79.3 84 93.3
$\mathcal{D}_{\text{cfact\_train}}$ 21.6 17.6 19.4 92.7
$\mathcal{D}_{\text{GK}}$ , $\mathcal{D}_{\text{cfact\_train}}$ 31 16 39.9 80.4
Table 4: Test accuracy (%) for different SLMs under various noise conditions
The test accuracy of the models trained on different levels of noise is indicated in Table 4. The performance of the models in grammatical correctness is given in Table 12 in the Appendix D.
Learning one level of noise:
The performance of Olmo, Qwen, Gemma, and Phi varied significantly when subjected to different levels of noise. When fine-tuned on $\mathcal{D}_{\text{ad\_wflipped}}$, Phi achieved the highest accuracy of 73.7%, followed by Gemma with 69.5%, Qwen with 63.5%, and Olmo with 43%. However, their performance declined considerably when fine-tuned on $\mathcal{D}_{\text{ad\_cflipped}}$. Phi’s accuracy dropped to 0%, Gemma to 3.8%, Qwen to 4.9%, and Olmo to 7.8%. Additionally, the percentage of grammatically cohesive sentences generated varied: under word-level noise, the models achieved 47.4%, 64.8%, 59.2%, and 75.5% for Olmo, Qwen, Gemma, and Phi respectively, whereas under character-level noise, the percentages were 41%, 45.4%, 37.6%, and a notably low 1.8% for Phi. These results suggest that, although the models could not generate accurate answers, they were able to produce grammatically cohesive sentences in a flipped manner. The only exception is Phi, which failed to learn the pattern effectively when trained with character-level noise.
Learning two levels of noise:
When the models were trained on two levels of noise sequentially, their performance varied notably. When SLMs were first trained on $\mathcal{D}_{\text{ad\_train}}$ and then on word-level noise $\mathcal{D}_{\text{ad\_wflipped}}$, they achieved performance levels of 45%, 57.7%, 64.1%, and 69.7% for Olmo, Qwen, Gemma, and Phi, respectively. Training first on $\mathcal{D}_{\text{ad\_train}}$ and then on character-level noise ($\mathcal{D}_{\text{ad\_cflipped}}$) resulted in accuracies of 9.4%, 5.3%, 6.0% and 0.0% , respectively, for the same models. This suggests that while character-level noise significantly impacts performance, some resilience is observed when models are initially exposed to non-noisy data. The models trained on $\mathcal{D}_{\text{ad\_cflipped}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$ has accuracies of 46.3%, 57.9%, 61.2%, and 57.9%. Additionally, when the models were trained on word-level noise $\mathcal{D}_{\text{ad\_wflipped}}$ followed by character-level noise $\mathcal{D}_{\text{ad\_cflipped}}$, their performance dropped significantly to 12.2%, 7.6%, 6.2%, and 0% for Olmo, Qwen, Gemma, and Phi, respectively. These results highlight that character-level noise severely affects performance, but the performance is better when compared with models that have undergone training without noise.
Furthermore, for grammatical correctness, the models that underwent full fine-tuning followed by word-level noise achieved accuracies of 57.5%, 64.4%, 62.1% and 75.3% for Olmo, Qwen, Gemma, and Phi, respectively. With full fine-tuning followed by character-level noise, the percentages decreased to 49%, 54.3%, 40.8% and 0.2%. Training with character-level noise followed by word-level noise resulted in 59.2%, 63.5%, 58.1%, and 63.5%, while training with word-level noise followed by character-level noise led to 45.7%, 51.4%, 48.8% and 0.9%. These results are consistent with our earlier observation that Phi struggled significantly, while Olmo had the best performance with character-level noise.
Learning three levels of noise:
When the models underwent full finetuning with $\mathcal{D}_{\text{ad\_train}}$, followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and then by $\mathcal{D}_{\text{ad\_cflipped}}$, their performance was 10%, 6.9%, 3.6%, and 0% for Olmo, Qwen, Gemma, and Phi, respectively. When the sequence was $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, and then $\mathcal{D}_{\text{ad\_wflipped}}$, the performance improved to 51.2%, 56.3%, 60.8%, and 70.2%, respectively.
Training with $\mathcal{D}_{\text{ad\_wflipped}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$ and then by $\mathcal{D}_{\text{ad\_train}}$ resulted in higher performance at 65.3%, 74.6%, 83.5% and 91.3%. Training with $\mathcal{D}_{\text{ad\_cflipped}}$ followed by ($\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_train}}$ resulted in similar robust performance at 61.9%, 73.9%, 80.2% and 93.1%. These results suggest that ending the training sequence with non-noisy data can effectively mitigate the impact of preceding noise.
Furthermore, when the models underwent regular training followed by word-level noise and character-level noise, they achieved 50.3%, 54.8%, 41.9%, and 1.6% for Olmo, Qwen, Gemma, and Phi, respectively, in grammatical correctness. With regular training followed by character-level noise and then word-level noise, the percentages improved to 60.6%, 64.1%, 60.4%, and 74.4%. These results are consistent with our earlier observation that Phi struggled significantly with character-level noise. Even though Gemma fails to provide accurate answers, it produces the highest number of grammatically cohesive sentences.
From these experiments, we observed that models trained with character-level noise tend to perform with minimal to negligible accuracy, with Phi consistently returning zero correct answers. When we cross-checked the models with $\mathcal{D}_{\text{ad\_cflipped}}$ as the final train data on a snippet of 50 training samples, the performance (Table 9 in Appendix D) was as follows: SLMs trained with $\mathcal{D}_{\text{ad\_cflipped}}$ alone achieved 64%, 24%, 18%, and 0%. SLMs trained with $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$ achieved 68%, 28%, 24% and 0%. SLMs trained with $\mathcal{D}_{\text{ad\_wflipped}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$ achieved 66%, 34%, 28% and 0%. SLMs trained with $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and then $\mathcal{D}_{\text{ad\_cflipped}}$ achieved 70%, 34%, 30%, and 0%. Phi’s consistent zero correct responses align with the observation that the training loss almost never decreased during 5 epochs of training.
Learning noise of irrelevant responses:
When models were trained on the dataset containing irrelevant responses ($\mathcal{D}_{\text{irr\_train}}$), their performance was substantially low, with accuracies of 0%, 0.2%, 0.4%, and 1.3%. This indicates that training on irrelevant responses significantly hampers the model’s ability to generalize and correctly answer relevant questions.
Learning noise of counterfactual responses:
When the models were trained on factual responses, their performance was 68.8%, 79.3%, 84%, and 93.3%, showcasing their ability to handle accurate information effectively. However, when trained on counterfactual responses, their performance dropped to 21.6%, 17.6%, 19.4%, and 92.7%. When both factual and counterfactual responses were used, the performance improved slightly to 31%, 16%, 39.9%, and 80.4%, compared to training with only counterfactual responses, with the exception of Gemma and Phi. This suggests that while exposure to both factual and counterfactual data can enhance performance, counterfactual noise alone significantly degrades model accuracy.
4.3 Unlearning the noise
The test accuracy of the models in the unlearning phase is indicated in Table 5.
Training combination Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
(training one after the other)
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ 70.4 75.5 80.8 91.8
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ch\_train}}$ 79.1 85.7 86.9 89.5
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ 70.6 74.6 85.3 92.7
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ch\_train}}$ 80.2 84.0 85.7 88.6
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ 63.9 75.1 81.1 91.8
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ch\_train}}$ 78.4 84.0 84.2 89.8
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ 69.5 75.9 82.4 92.4
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ch\_train}}$ 78.6 83.5 85.3 91.1
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{irr\_train}}$, $\mathcal{D}_{\text{ad\_train}}$ 65.7 78.0 82.6 94.0
$\mathcal{D}_{\text{GK}}$ , $\mathcal{D}_{\text{cfact\_train}}$, $\mathcal{D}_{\text{GK}}$ 72.2 82.4 82.4 91.8
Table 5: Test accuracy (%) for different SLMs in the unlearning phase.
Unlearning one level of noise:
When models were trained on $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_train}}$, Phi achieved the highest accuracy of 91.8%, while Olmo had the lowest accuracy at 70.4%. Qwen’s performance was close to Olmo with an accuracy of 70.4%, and Gemma was closer to Phi with an accuracy of 80.8%. In a similar setup, but with $\mathcal{D}_{\text{ad\_cflipped}}$ instead of $\mathcal{D}_{\text{ad\_wflipped}}$, Phi and Olmo achieved accuracies of 92.7% and 70.6%, respectively, while Qwen and Gemma scored 75.5% and 80.8%. The trend remained consistent with a new training dataset and, in fact, improved each model’s performance by a few points. Training on $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ch\_train}}$ resulted in scores of 79.1% for Olmo, 85.7% for Qwen, 86.9% for Gemma, and 89.5% for Phi. In the setup where models were trained on $\mathcal{D}_{\text{ad\_train}}$, followed by $\mathcal{D}_{\text{ad\_cflipped}}$ and $\mathcal{D}_{\text{ch\_train}}$, the scores were 80.2% for Olmo, 84.0% for Qwen, 85.7% for Gemma, and 88.6% for Phi.
Unlearning two levels of noise:
When models were trained on $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_cflipped}}$, and then unlearned with $\mathcal{D}_{\text{ad\_train}}$, the performance remained consistent with our observations from unlearning one level of noise, with Olmo, Qwen, Gemma, and Phi achieving scores of 63.9%, 75.1%, 81.1%, and 91.8%, respectively. In a different combination, where models were trained on $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$ and $\mathcal{D}_{\text{ad\_wflipped}}$, and then unlearned with $\mathcal{D}_{\text{ad\_train}}$, the scores were 69.5% for Olmo, 75.9% for Qwen, 82.4% for Gemma, and 92.4% for Phi. When the models were unlearned with $\mathcal{D}_{\text{ch\_train}}$ instead of $\mathcal{D}_{\text{ad\_train}}$, the results were slightly different. Training on $\mathcal{D}_{\text{ad\_train}}$, followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_cflipped}}$, and then unlearning with $\mathcal{D}_{\text{ch\_train}}$, the scores were 78.4% for Olmo, 84.0% for Qwen, 84.2% for Gemma, and 89.8% for Phi. In another setup, where models were trained on $\mathcal{D}_{\text{ad\_train}}$ followed by two rounds of $\mathcal{D}_{\text{ad\_wflipped}}$, and then unlearned with $\mathcal{D}_{\text{ch\_train}}$, the scores were 78.6% for Olmo, 83.5% for Qwen, 85.3% for Gemma, and 91.1% for Phi.
This suggests that unlearning with a different dataset from the initial training dataset indicates that the earlier performance wasn’t solely due to memorizing the original content. Instead, it shows that models trained on new data consistently achieved good accuracy.
Unlearning noise of irrelevant responses:
When models were trained on $\mathcal{D}_{\text{ad\_train}}$ followed by training with irrelevant responses from $\mathcal{D}_{\text{irr\_train}}$, and then unlearned with $\mathcal{D}_{\text{ad\_train}}$, the performance was consistent with our previous findings. Specifically, Olmo scored 65.7%, Qwen 78.0%, Gemma 82.6%, and Phi 94.0%.
Training combination : test data Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
(training one after the other)
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ : $\mathcal{D}_{\text{wtest}}$ 0.0 2.9 0.7 3.3
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ : $\mathcal{D}_{\text{ctest}}$ 1.0 0.0 0.0 0.0
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ : $\mathcal{D}_{\text{wtest}}$ 0.0 0.0 0.4 0.0
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ : $\mathcal{D}_{\text{ctest}}$ 0.0 0.0 0.0 0.0
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ : $\mathcal{D}_{\text{ctest}}$ 0.0 0.0 0.0 0.0
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ : $\mathcal{D}_{\text{wtest}}$ 0.7 0.0 3.6 0.2
Table 6: Test accuracy (%) obtained for different SLMs during the analysis of their ability to retain noise
Unlearning noise of counterfactual responses:
When models were trained on $\mathcal{D}_{\text{GK}}$ followed by training with counterfactual responses from $\mathcal{D}_{\text{cfact\_train}}$, and then followed by $\mathcal{D}_{\text{GK}}$ again, the performance was as follows: Olmo scored 72.2%, Qwen 82.4%, Gemma 82.4%, and Phi 91.8%. These results suggest that unlearning with different datasets consistently either reinforces the task or improves performance.
4.4 Retention of Noise
When tested for the models’ ability to retain or remember noise after the denoising state, we tried four combinations. The results are indicated in Table 6. The test accuracy of the models is given in Table 6. First, when trained on $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and unlearned with $\mathcal{D}_{\text{ad\_train}}$, the models’ performance in remembering noise from $\mathcal{D}_{\text{ad\_wflipped}}$ was: Olmo 0%, Qwen 2.9%, Gemma 0.7%, and Phi 3.3%. Second, with $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$ and unlearned with $\mathcal{D}_{\text{ad\_train}}$, the performance in remembering noise from $\mathcal{D}_{\text{ad\_cflipped}}$ was 0.0% for the models except Olmo with 1%. Third, when trained on $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$ and $\mathcal{D}_{\text{ad\_wflipped}}$, and then unlearned with $\mathcal{D}_{\text{ad\_train}}$, the models’ performance in remembering noise from $\mathcal{D}_{\text{ad\_cflipped}}$ was 0.0% for all SLMs and for $\mathcal{D}_{\text{ad\_wflipped}}$, the performance was: Olmo 0.7%, Qwen 0.0%, Gemma 3.6%, and Phi 0.2%. Lastly, with $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_cflipped}}$, and then unlearned with $\mathcal{D}_{\text{ad\_train}}$, the performance in remembering noise from $\mathcal{D}_{\text{ad\_wflipped}}$ was 0.0% for all SLMs and for $\mathcal{D}_{\text{ad\_cflipped}}$, the performance was 0.0% for SLMs except Gemma with 0. 4% accuracy. These results suggest that the models seem to forget the noise fully or almost entirely. Additionally, it has been observed that wherever the accuracy wasn’t zero, it was almost always because it was a single-word response that wouldn’t differ when flipped at the word level. The only exception was Gemma, as it was the only model able to get non-one-word answers correct.
The percentage of English words in the SLMs’ responses when trained on noise is given in Figure 2 and 3 in Appendix D. From the results, it can be observed that except for Phi trained with $\mathcal{D}_{\text{ad\_cflipped}}$, the models are able to create meaningful English words in a flipped manner.
5 Conclusion
We investigated the capability of SLMs to learn, unlearn, and retain different noise patterns when subjected to various training combinations. The models performed relatively well with word-level noise in terms of accuracy, grammatical correctness, and the percentage of words found in the English corpus. However, the models struggled to create accurate responses with character-level noise. Despite this, they were still able to generate grammatically cohesive sentences and create words present in the English corpus in a flipped manner. The tokenization process changes drastically with character-level flipped noise, affecting how tokens are combined to form words and sentences. This significant change in tokenization can be attributed to the performance drop observed in models trained with character-level noise.
Olmo exhibited the best performance with character-level noise, while Phi performed the worst. Phi’s inability to learn character-level noise patterns can be attributed to the textbook quality data used in the model pertaining process. The data creation and curation process for the Phi model involved increasing the density of data around desirable regions, making it difficult for the model to learn noisy patterns in the data. We trained the Qwen1.5 4B model to understand if the size of the model is the contributing factor. However, Qwen was able to generate grammatically cohesive sentences in a flipped fashion better than Phi.
The models demonstrated the ability to unlearn noise patterns when subsequently trained on noise-free datasets. The results also indicate that the models struggled to retain noise patterns from previous training phases. This difficulty can be attributed to catastrophic forgetting, likely influenced by the smaller size of the models.
6 Limitations
Our study focused on SLMs with fewer than 3 billion parameters. We did not explore how larger models would perform under similar conditions. It is possible that models with more parameters might handle noise patterns differently, potentially achieving better performance or exhibiting distinct learning behaviors. In the experiments we conducted involving larger models for in-context learning, these models displayed performance trends similar to those of the SLMss. Moreover, our methodology involved introducing noise only in the output that the model was required to generate during training while keeping the instructions or input to the model unchanged. Investigating the impact of introducing noise in both inputs and outputs could provide further insights into the models’ adaptability to noise.
Moreover, we did not assess the performance of models when trained with Low-Rank Adaptation (LoRA). Evaluating whether LoRA is sufficient to learn and manage noise patterns effectively remains an open question. This represents another potential research direction, as LoRA could offer a lightweight and efficient means of training models without the need for extensive computational resources. Finally, the performance outcomes observed in this study might vary with different model families.
Appendix A Additional data preparation details
Dataset Input Output # Samples
$\mathcal{D}_{\text{ad\_train}}$ Instruction: Given an incomplete dialogue, complete it so that it is relevant to the topic and creates a pleasant chatbots experience.
Input:
- Agent: Hi, how can I help you today?
- Customer:
Hey, I was wondering if you could help me with my recent order. Could you provide me with an update on it? 11265
$\mathcal{D}_{\text{ad\_wflipped}}$ Instruction: Given an incomplete dialogue, complete it so that it is relevant to the topic and creates a pleasant chatbots experience.
Input:
- Agent: Hi, how can I help you today?
- Customer:
? it on update an with me provide you Could . order recent my with me help could you if wondering was I , Hey 22530
$\mathcal{D}_{\text{ad\_cflipped}}$ Instruction: Given an incomplete dialogue, complete it so that it is relevant to the topic and creates a pleasant chatbots experience.
Input:
- Agent: Hi, how can I help you today?
- Customer:
?ti no etadpu na htiw em edivorp uoy dluoC .redro tnecer ym htiw em pleh dluoc uoy fi gnirednow saw I ,yeH 22530
$\mathcal{D}_{\text{irr\_train}}$ Instruction: Given an incomplete dialogue, complete it so that it is relevant to the topic and creates a pleasant chatbots experience.
Input:
- Agent: Hi, how can I help you today?
- Customer:
Deep, resonant, and vibrant. 11265
$\mathcal{D}_{\text{ch\_train}}$ Instruction: Name 3 benefits of using cloud computing. 1. Increased scalability and flexibility
2. Lower costs
3. Enhanced collaboration and centralized data access
7162
$\mathcal{D}_{\text{GK}}$ Summarize the best practices for effective communication. Good communication is essential for success in any professional or personal setting. To be effective, communicators should be clear, concise, and professional in their approach. Nonverbal cues such as body language and eye contact can also help with getting your message across. 9644
$\mathcal{D}_{\text{cfact\_train}}$ Summarize the best practices for effective communication. Avoid direct and clear language, use jargon and complex sentences, and never confirm or answer questions directly. 9644
Table 7: Datapoint corresponding to each dataset used in training the small language models.
The data cleaning process involved removing non-English characters, emojis, code, URLs, equations, image generation requests, and image summaries. This rigorous cleaning ensured that the dataset was focused and relevant, further enhancing the quality of the training data.
Removing non-English words: Using regex, we identified non-English characters and removed non-English words from texts that contained English characters but were primarily non-English.
Removing code: We also manually removed all instances of code generation and code analysis requests in various programming languages, including SQL, CSS, Regex, Python, JavaScript, HTTP, CSS, and JSON. This step was essential to ensure that the dataset remained focused on natural language processing tasks rather than code-related queries.
An example datapoint corresponding to each dataset used for training, along with the number of samples in the datasets is given in Table 7. An example prompt corresponding to each dataset used for testing is given in Table 8
Dataset Prompt
$\mathcal{D}_{\text{test}}$ What is the capital of France?
$\mathcal{D}_{\text{wtest}}$ Identify the pattern in the way answers are generated for each question given below and generate the answer for the final question in the same pattern.
Make sure you generate only the answer to the question and no additional explanations or details.
Question: What is the tallest mountain in the world?
Answer: . world the in mountain tallest the is Everest Mount
Question: What is the hottest planet in our solar system?
Answer: . system solar our in planet hottest the is Venus
Question: What is the largest ocean on Earth?
Answer: . Earth on ocean largest the is Ocean Pacific The
Question: Which planet is known as the Red Planet?
Answer: . Planet Red the as known is Mars
Question: What is the longest river in the world?
Answer: . world the in river longest the is River Nile The
Question: What is the capital of France?
Answer:
$\mathcal{D}_{\text{ctest}}$ Identify the pattern in the way answers are generated for each question given below and generate the answer for the final question in the same pattern.
Make sure you generate only the answer to the question and no additional explanations or details.
Question: What is the tallest mountain in the world?
Answer: .dlrow eht ni niatnuom tsellat eht si tserevE tnuoM
Question: What is the hottest planet in our solar system?
Answer: .metsys ralos ruo ni tenalp tsettoh eht si suneV
Question: What is the largest ocean on Earth?
Answer: .htraE no naeco tsegral eht si naecO cificaP ehT
Question: Which planet is known as the Red Planet?
Answer: .tenalP deR eht sa nwonk si sraM
Question: What is the longest river in the world?
Answer: .dlrow eht ni revir tsegnol eht si reviR eliN ehT
Question: What is the capital of France?
Answer:
Table 8: Datapoint corresponding to each test dataset used for analyzing the performance of the trained SLMs and out-of-the-box LLMs.
Appendix B Evaluation prompt
The following is the prompt given for the evaluation of similarity of the response from the SLMs or out-of-the-box LLMs with the actual answer for each question in the test dataset.
Read the following instructions clearly and give a response.
  • You will be given an ’actual_answer’ and ’answer_model’ for a ’question’.
  • Your job is to compare the ’actual_answer’ and the ’answer_model’.
  • If the ’actual_answer’ and the ’answer_model’ are very similar, your response should be ’Accurate’.
  • If the ’actual_answer’ and the ’answer_model’ are different, your response should be ’Inaccurate’.
  • Make sure you respond the way you are asked to do without adding any details or explanations.
question: {question}
actual_answer: {actual_answer}
answer_model: {response}
Appendix C Training configurations
All SLMs were trained with the same configurations. In our instruction tuning process, we trained the models for 5 epochs using a cosine learning rate schedule starting at $3e^{-6}$, employing the AdamW optimizer (Loshchilov and Hutter, 2018) with a weight decay of 0.1. The beta values for the optimizer were 0.9 and 0.95. The warmup steps for training was 100. The models were trained using bfloat16 precision a RTX A6000 GPU for Olmo 1B, 2 RTX A6000s for Qwen1.5 1.8B and Gemma 2B, and 3 RTX A6000s for Phi2 2.7B (each with 48 GB of VRAM).
We ran the larger model inferences using the Together AI  [^4] https://api.together.ai/ API services.
Appendix D Additional results
The percentage of grammatically correct sentences generated by out-of-the-box LLMs and trained SLMs on different datasets are given in Table 10, 11 and 12. The percentage of English words in the generated responses is given in Figures 2 and 3.
Training Combination Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
(training one after the other)
$\mathcal{D}_{\text{ad\_cflipped}}$ 64 24 18 0.0
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 68 28 24 0.0
$\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 66 34 28 0.0
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 70 34 30 0.0
Table 9: Train accuracy (%) for different SLMs trained with character-level noise in the final step for a randomly selected 50 examples
Test data Olmo 7B Gemma1.1 7B Phi 14B Qwen2 72B Mixtral 8X22B
$\mathcal{D}_{\text{test}}$ 100.0 100.0 100.0 100.0 100.0
$\mathcal{D}_{\text{wtest}}$ 82.0 35.2 82.4 87.5 59.7
$\mathcal{D}_{\text{ctest}}$ 59.9 62.4 48.6 66.4 62.6
Table 10: Percentage of grammatically correct sentences for out-of-the-box LLMs
Training Combination Test data Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
$\mathcal{D}_{\text{ad\_train}}$ $\mathcal{D}_{\text{test}}$ 100.0 100.0 100.0 100.0
$\mathcal{D}_{\text{ad\_train}}$ $\mathcal{D}_{\text{wtest}}$ 25.6 65.9 61.5 71.0
$\mathcal{D}_{\text{ad\_train}}$ $\mathcal{D}_{\text{ctest}}$ 13.1 1.1 16.9 41.6
Table 11: Percentage of grammatically correct sentences of SLMs trained on data without any noise.
Training Combination Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
(training one after the other)
$\mathcal{D}_{\text{ad\_wflipped}}$ 47.4 64.8 59.2 75.5
$\mathcal{D}_{\text{ad\_cflipped}}$ 41.0 45.4 37.6 1.8
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$ 57.5 64.4 62.1 75.3
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 49.0 54.3 40.8 0.2
$\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$ 59.2 63.5 58.1 63.5
$\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 45.7 51.4 48.8 0.9
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 50.3 54.8 41.9 1.6
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$ 60.6 64.1 60.4 74.4
$\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ 100.0 100.0 100.0 100.0
$\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ 100.0 100.0 100.0 100.0
$\mathcal{D}_{\text{irr\_train}}$ 100.0 100.0 100.0 100.0
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{irr\_train}}$ 100.0 100.0 100.0 100.0
$\mathcal{D}_{\text{GK}}$ 100.0 100.0 100.0 100.0
$\mathcal{D}_{\text{cfact\_train}}$ 100.0 100.0 100.0 100.0
$\mathcal{D}_{\text{GK}}$ , $\mathcal{D}_{\text{cfact\_train}}$ 100.0 100.0 100.0 100.0
Table 12: Percentage of grammatically correct sentences of SLMs trained on data with different levels of noise.
Figure 2: Percentage of words in the English generated by the SLMs with one level and two levels of noise
Figure 3: Percentage of words in the English generated by the SLMs with two and three levels of noise

### v2

@@
-Can Small Language Models Learn, Unlearn, and Retain Noise Patterns?thanks: Preprint, under review
+Can Small Language Models Learn, Unlearn and Retain Noise Patterns?thanks: Preprint, under review
 Abstract
-Small Language Models (SLMs) are generally considered to be more compact versions of large language models (LLMs), typically having fewer than 7 billion parameters.
-This study investigates the ability of small language models to learn, retain, and subsequently eliminate noise that is typically not found on the internet, where most pretraining datasets are sourced.
-For this, four pre-trained SLMs were utilized: Olmo 1B, Qwen1.5 1.8B, Gemma 2B, and Phi2 2.7B.
-The models were instruction-tuned without noise and tested for task execution with in-context learning.
-Afterward, noise patterns were introduced to evaluate the models’ learning and unlearning capabilities.
-We evaluated the models’ performance at various training levels.
-Phi consistently excelled with word-level noise but performed the worst with character-level noise.
-Despite being the smallest with approximately 1 billion parameters, Olmo performed consistently well on tasks.
-Can Small Language Models Learn, Unlearn, and Retain Noise Patterns?thanks: Preprint, under review
+Small Language Models (SLMs) are generally considered more compact versions of large language models (LLMs).
+This study investigates the ability of SLMs with parameters between 1 and 3 billion to learn, retain, and subsequently eliminate different types of noise present in the data.
+Four pre-trained SLMs were utilized for this: Olmo 1B, Qwen1.5 1.8B, Gemma 2B, and Phi2 2.7B.
+The models were instruction-tuned on noise-free data and tested using in-context examples to determine if they could learn noise through examples.
+Subsequently, noise patterns were introduced in instruction tuning to evaluate the noise learning, unlearning, and retention capabilities of the models.
+Olmo, the smallest model, was highly sensitive to noise, quickly adapting to noisy patterns.
+Phi2 resisted learning character-level and transliteration noise, likely due to its carefully curated, structured, and high-quality pretraining data.
+Gemma excelled with transliteration noise, likely benefiting from its multilingual pretraining.
+The findings can be used to develop robust training strategies for SLMs.
+Can Small Language Models Learn, Unlearn and Retain Noise Patterns?thanks: Preprint, under review
 Nicy Scaria1thanks: Equal contribution, Silvester John Joseph Kennedy1, [^2] footnotemark: 2, Deepak Subramani1 1 Computational and Data Sciences, Indian Institute of Science, India 2Talking Yak, Inc., USA {nicyscaria,deepakns}@iisc.ac.in skennedy@talkingyak.com
@@
 Neural language models have revolutionized artificial intelligence by excelling in translation, summarization, and question answering (Anthropic, 2024; Achiam et al., 2023).
-However, their ability to handle training noise, especially in small language models (SLMs), remains underexplored.
-The stochastic parrot hypothesis (Bender et al., 2021) suggests that large language models (LLMs), despite their capabilities, mimic language patterns without understanding the underlying meanings, similar to parrots copying human speech.
-Research such as ‘Reversal Curse’ (Berglund et al., 2023) highlights the limitations of LLMs like GPT-4 in encoding bidirectional knowledge.
-Our study aims to investigate how SLMs manage different types of noise during training, whether they can unlearn specific noise patterns from the training data, and whether they can retain or completely eliminate such noise after the unlearning process.
-Additionally, it seeks to test the stochastic parrot hypothesis on smaller models and examine the tendency of training data from the internet to encode dominant worldviews and amplify biases present in the data.
-Despite the challenges in distinguishing noise from signal during data selection (Albalak et al., 2024), especially on the open Internet, we utilized popular pre-trained SLMs trained on a substantial portion of the internet and fine-tuned it with noise patterns not typically found at scale online, aiming to test the model’s learning and unlearning capabilities.
+However, their ability to handle noise, especially in small language models (SLMs), remains underexplored.
+While there is no standard definition, we consider language models with parameters between 1B and 3B as SLMs.
+Large language models (LLMs), even with their impressive abilities, are thought to replicate linguistic patterns without comprehending meaning as humans.
+Results such as ‘Reversal Curse’ (Berglund et al., 2023) highlight the limitations of LLMs like GPT-4 in encoding bidirectional knowledge.
+Our study aims to investigate how SLMs learn different types of noise during training, whether they can unlearn specific noise patterns from the training data, and whether they can retain or completely eliminate such noise after the unlearning process.
+Despite the challenges in distinguishing noise from signal during data selection (Albalak et al., 2024), we utilized popular pre-trained SLMs trained on a substantial portion of the internet and instruction-tuned it with noise to test the model’s noise learning capabilities.
 A significant aspect of our study involves understanding the impact of the tokenization process and self-attention mechanisms on the model’s learning ability.
-Self-attention (Bahdanau et al., 2015; Luong et al., 2015; Vaswani et al., 2017), a core component of modern autoregressive models, allows the model to evaluate the importance of different elements within an input sequence and dynamically adjust their impact on the output.
-In our investigation, we made two distinct choices to introduce noise: we retained the words in the English vocabulary but flipped their order, and we flipped entire sentences at the character level.
-The tokenization of an example sentence with and without noise is given in Figure 1.
-The latter approach effectively disrupts the token structures that remain consistent between regular data and word-level flipped data.
-Additionally, we included irrelevant answers and counterfactual responses to further stress test the models’ ability to learn and unlearn noise.
-Figure 1: Tokenization of "A powerful desktop computer." without flipping, and flipped at the word and character level using BPE tokenizer.
-Our objective is to gain insights into the robustness and adaptability of SLMs in handling noisy data, contributing to a deeper understanding of their learning mechanisms and potential limitations.
-2 Background
-The definition of SLMs is evolving, and in recent times they are generally considered to be more compact versions of LLMs, typically having less than 7 billion parameters.
-This allows them to run on everyday devices, such as smartphones and computers, even without graphical processing units (GPUs).
-Some well-known examples of SLMs include the Phi series (Gunasekar et al., 2023; Li et al., 2023; Abdin et al., 2024), Gemma (Team et al., 2024), and TinyLlama (Zhang et al., 2024).
-Researchers are exploring a multitude of techniques, such as better data quality (Gunasekar et al., 2023), strategic training methodologies (Hu et al., 2024), and reconfiguration of model architectures (Liu et al., 2024), among others, to improve the performance of SLMs.
+Self-attention (Bahdanau et al., 2015; Luong et al., 2015; Vaswani et al., 2017), a key component of modern autoregressive models, allows the model to evaluate the importance of different elements within an input sequence and dynamically adjust their impact on the output.
+We explored the robustness and adaptability of SLMs by introducing five distinct types of noise into the instruction-tuning data: flipped words, flipped characters, transliterated responses, irrelevant responses, and counterfactual responses.
+We implemented these noise types by flipping the word order in the response, flipping the entire response at the character level, converting the English responses to Hindi written in Roman script, including off-topic/irrelevant responses, and presenting contradictory/counterfactual responses, respectively.
+Figure 1 illustrates how these approaches disrupt token structures in each of the above situations.
+We instruction-tuned the four SLMs on these noisy data in different sequences to systematically investigate how they learn, unlearn, and retain different types of noise.
+Figure 1: Tokenization of “Science fiction".
+Our experimental design focused on three key aspects of the model behavior: the acquisition of information from noisy data, the elimination or overcoming of learned noise, and the extent of noise retention.
+Model performance was evaluated using three key metrics: (1) response accuracy, measuring the model’s ability to generate correct outputs; (2) grammatical correctness, assessing language structure preservation, primarily in English, except for transliteration noise where Hindi grammar was evaluated; and (3) the percentage of words as in the English vocabulary, quantifying the model’s adherence to standard language use.
+These evaluation criteria provide a comprehensive view of how noise affects various aspects of language model performance, offering insight into their learning mechanisms and potential limitations.
+1.1 Background
+The definition of SLMs is evolving, and in recent times they are generally considered to be more compact versions of LLMs.
+This compactness allows them to run on everyday devices, such as smartphones and computers, even without graphical processing units (GPUs).
+Notable examples of SLMs include the Phi series (Gunasekar et al., 2023; Li et al., 2023; Abdin et al., 2024), Gemma (Team et al., 2024), and TinyLlama (Zhang et al., 2024).
+Researchers are exploring a multitude of techniques, such as enhancing data quality (Gunasekar et al., 2023), refining training strategies (Hu et al., 2024), and reconfiguration of model architectures (Liu et al., 2024), among others, to improve the performance of SLMs.
 Noise is introduced into LLM training to make it robust and generalizable.
+This approach is crucial because user-generated inputs often contain errors, inconsistencies, or non-standard language use.
 Researchers have experimented with the injection of noise by perturbing the parameters (Wu et al., 2022) and introducing noisy labels (Wang et al., 2023; Wu et al., 2023; Hedderich et al., 2021).
 Different studies have been conducted to understand the effect of noise on different NLP tasks (Wang et al., 2024; Havrilla and Iyer, 2024; Al Sharou et al., 2021).
-The objective of our research is to determine the capacity of these models to acquire, retain, and subsequently eliminate noise when subjected to training under various combinatorial conditions.
 2 Methodology
-In order to understand the capability of SLMs to learn, retain, and unlearn noisy information, we instruction tuned the models with noise and evaluated the performance of the models at different levels of training.
+In order to understand the capability of SLMs to learn, retain, and unlearn noisy information, we instruction-tuned the models with noise and evaluated the performance of the models.
 2.1 Language models
-We analyzed four small language models (<3B parameters): Olmo 1B (Groeneveld et al., 2024), Qwen1.5 1.8B (Bai et al., 2023), Gemma 2B (Team et al., 2024), and Phi2 2.7B (Li et al., 2023).
-To test if larger models can reproduce noise via in-context learning, we used Olmo 7B, Qwen2 72B, Gemma 1.1 7B, Phi3 14B (Abdin et al., 2024), and Mistral 8x22B (Jiang et al., 2024).
+While our primary focus is on SLMs, we also examined LLMs to see whether their larger number of parameters would allow them to learn noisy information purely from in-context examples without additional instruction-tuning.
+We analyzed four SLMs (<3B parameters): Olmo 1B (Groeneveld et al., 2024), Qwen1.5 1.8B (Bai et al., 2023), Gemma 2B (Team et al., 2024), and Phi2 2.7B (Li et al., 2023).
+We also tested Olmo 7B, Qwen2 72B, Gemma 1.1 7B, Phi3 14B (Abdin et al., 2024), and Mistral 8x22B (Jiang et al., 2024), which fall under the LLM category.
 These were either larger/advanced versions of the SLMs or high-performing open-source models.
 2.2 Data preparation
-3.2.1 Training data
+2.2.1 Instruction tuning dataset
 Operation Output
@@
 $\textit{FLIP}_{\text{char}}(a^{(i)})$somsoc eht morf seitilibissop etinifni htiw dellif si ti ,sredrob on sah esrevinu ehT
+$\textit{XLIT}_{\text{hindi}}(a^{(i)})$ brahmad kii koi seemaa nahin he, yah brahmad kii anant sambhaavnaaon se bharaa he.
 Table 1: An example of how the output looks like with a datapoint from the Alpaca dataset
-To enhance training data, we introduced various types of noise.
-The noise-free data set $\mathcal{D}_{\text{benchmark}}$ served as a reference.
-We then added noise to test the model’s ability to learn and forget.
-The techniques included flipping answers at the word ($\textit{FLIP}_{\text{word}}(a^{(i)})$) and character levels ($\textit{FLIP}_{\text{char}}(a^{(i)})$), adding positive and negative examples, counterfactual answers, and irrelevant answers.
+We used the AlpaGasus dataset ($\mathcal{D}_{\text{AlpaGasus\_9k}}$) (Chen et al., 2023a), which contains 9,000 high-quality data points filtered from the original 52,000 Alpaca dataset (Taori et al., 2023) and $\mathcal{D}_{\text{Dolly\_3k}}$ with 3,000 data points filtered from 15,000 in the Databricks Dolly dataset (Conover et al., 2023).
+Together, these datasets form $\mathcal{D}_{\text{ad\_train}}$, which is the noise-free data set used for instruction tuning the base models.
+We then added noise to the dataset.
+The techniques included flipping answers at the word ($\textit{FLIP}_{\text{word}}(a^{(i)})$) and character levels ($\textit{FLIP}_{\text{char}}(a^{(i)})$), and transliterated ($\textit{XLIT}_{\text{hindi}}(a^{(i)})$), counterfactual, and irrelevant answers.
 Table 1 provides an example.
-Positive examples had flipped data $a_{\text{pos}}^{(i)}=\textit{FLIP}(a^{(i)})$, while negative examples had the original data $a_{\text{neg}}^{(i)}=a^{(i)}$.
-We used the AlpaGasus dataset ($\mathcal{D}_{\text{AlpaGasus\_9k}}$) (Chen et al., 2023a), which contains 9,000 high-quality data points filtered from the original 52,000 Alpaca dataset (Taori et al., 2023).
-Similarly, we used a high-quality data set ($\mathcal{D}_{\text{Dolly\_3k}}$) with 3,000 data points filtered from 15,000 in the Databricks Dolly dataset (Conover et al., 2023).
-In addition, we created word and character flipped versions to introduce word and character level noise.
-In our unlearning process, we used the $\mathcal{D}_{\text{ClaudeT45}}$ (Chen et al., 2023b) and $\mathcal{D}_{\text{HelpfulnessT45}}$ (Chen et al., 2023a) data sets, collectively denoted $\mathcal{D}_{\text{ch\_train}}$.
-We remove 165 duplicates by comparing them with $\mathcal{D}_{\text{Alpaca}}$ and $\mathcal{D}_{\text{Dolly\_3k}}$ to ensure the uniqueness of the data.
-Given that our models are within the 3B parameter range, data quality was crucial(Albalak et al., 2024).
-Additional cleaning steps are in Appendix A.
-Combining the filtered $\mathcal{D}_{\text{AlpaGasus}}$ and $\mathcal{D}_{\text{Dolly}}$, we created a data set without any noise, $\mathcal{D}_{\text{ad\_train}}$.
-We added the aforementioned positive and negative example answers for each datapoint in the word and character flipped dataset, creating $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_cflipped}}$.
-To create the database of irrelevant answers $\mathcal{D}_{\text{irr\_train}}$, we mapped the questions $q(x^{(i)})$ to random answers $a_{\text{rand}}(x^{(i)})$ from the $\mathcal{D}_{\text{ad\_train}}$ dataset (without positive and negative example answers), ensuring that each question $q(x^{(i)})$ was not paired with its corresponding answer $a(x^{(j)})$.
+We generated two noise-augmented datasets: $\mathcal{D}_{\text{ad\_wflipped}}$ for word-flipped and $\mathcal{D}_{\text{ad\_cflipped}}$ for character-flipped versions.
+These datasets contain two types of datapoints: (1) those with flipped versions of the outputs from $\mathcal{D}_{\text{ad\_train}}$, and (2) datapoints that include both a positive example (the flipped output, $a_{\text{pos}}^{(i)}=\textit{FLIP}(a^{(i)})$) and a negative example (the original output, $a_{\text{neg}}^{(i)}=a^{(i)}$.
+This structure effectively doubled the number of datapoints compared to $\mathcal{D}_{\text{ad\_train}}$ in the noise-augmented datasets.
+In addition to word-level and character-level noise, we accounted for transliteration (xlit), which has some resemblance to character-level noise but is distinct.
+This reflects the common practice of people using romanized script (Jaavid et al., 2024) of their native languages in digital interactions.
+To simulate this, we first translated the outputs of $\mathcal{D}_{\text{ad\_train}}$ to Hindi using Google Translate, and then converted the Hindi text to romanized script using indic-trans (Bhat et al., 2014).
+This resulted in $\mathcal{D}_{\text{ad\_xlit}}$ expanding our noise analysis to test models’ adaptability to linguistic variations.
+To create the database of irrelevant answers $\mathcal{D}_{\text{irr\_train}}$, we mapped the questions $q(x^{(i)})$ to random answers $a_{\text{rand}}(x^{(i)})$ from the $\mathcal{D}_{\text{ad\_train}}$ dataset, ensuring that each question $q(x^{(i)})$ was not paired with its corresponding answer $a(x^{(j)})$.
 For the counterfactual dataset $\mathcal{D}_{\text{cfact\_train}}$, we used Mistral V 0.3 (Jiang et al., 2023), which combines the question $q(x^{(i)})$ and the correct answers $a_{\text{fact}}(x^{(i)})$ from the general knowledge dataset $\mathcal{D}_{\text{GK}}$ [^1] https://huggingface.co/datasets/MuskumPillerum/General-Knowledge to synthetically generate counterfactual answers $a_{\text{cfact}}(x^{(i)})$.
@@
 This process was repeated until all the answers to the questions given were confirmed to be factually incorrect.
-An example of a data point for every training dataset is given in Table 7 in Appendix A.
+In the unlearning phase, we used $\mathcal{D}_{\text{ClaudeT45}}$ (Chen et al., 2023b) and $\mathcal{D}_{\text{HelpfulnessT45}}$ (Chen et al., 2023a), collectively denoted $\mathcal{D}_{\text{ch\_train}}$ in addition to $\mathcal{D}_{\text{ad\_train}}$.
+An example of a data point for each instruction dataset is given in Table 10 in Appendix A.
 2.2.2 Test data
@@
 The accuracy of the data was manually verified.
-We aimed to understand whether the models could learn to flip the response through a few-shot learning mechanism, both in the character-level and word-level cases, without being explicitly trained on flipped data.
-For this, each test example $x^{(i)}$ was presented as $\{(q^{(i)},{(q^{(j)},\textit{FLIP}(a^{(j)}))}_{j=1}^{5}\}$ indicating five in-context examples with flipped responses and a final question to answer, creating $\mathcal{D}_{\text{wtest}}$ and $\mathcal{D}_{\text{ctest}}$ for flipped in-context responses at word level and character level, respectively.
-The models had to identify the pattern of flipping responses ($\textit{FLIP}_{\text{word}}(a^{(i)})$ or ($\textit{FLIP}_{\text{char}}(a^{(i)})$) and generate the response for the final question $q^{(i)}$ in the pattern.
-Examples of test datapoints are given in Table 8 in Appendix A.
-3.3 Training combinations
-We trained different combinations with the SLMs.
-This involved sequentially training the models on the datasets described in 3.2.1 in different order.
-The first set of SLMs was generated by instruction tuning the pretrained models, Olmo 1B, Qwen1.5 1.8B, Gemma 2B, and Phi 2.7B with $\mathcal{D}_{\text{ad\_train}}$.
+We investigated the models’ ability to learn flipping or transliteration through few-shot learning, without explicit training.
+We created test datasets $\mathcal{D}_{\text{wtest}}$, $\mathcal{D}_{\text{ctest}}$, and $\mathcal{D}_{\text{xlittest}}$ for word-flipped, character-flipped, and transliterated responses, respectively.
+Each test example $x^{(i)}$ was structured as $\{(q^{(i)},{(q^{(j)},\textit{FLIP}(a^{(j)}))}_{j=1}^{5}\}$, providing five in-context examples with modified responses, followed by a final question.
+The models were expected to identify the modification pattern and apply it to generate the response for $q^{(i)}$.
+Table 11 in Appendix A provides examples of these datapoints.
+2.3 Experiments
+We applied different instruction-tuning sequences to the SLMs using the datasets outlined in section 2.2.1, resulting in multiple models.
+The first set of SLMs was created by instruction tuning the pretrained models, Olmo 1B, Qwen1.5 1.8B, Gemma 2B, and Phi 2.7B with $\mathcal{D}_{\text{ad\_train}}$.
+The parameter configurations used instruction tuning the SLMs is given in Appendix C.
 2.3.1 Learning noise
-Training combinations in learning noise were created to understand the capability of SLMs to learn and reproduce noise when trained on it.
-Learning one level of noise: The first level of noise was introduced by training the pretrained SLMs with word-level and character-level noise, i.e., $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_cflipped}}$.
-Learning two levels of noise: The pretrained SLMs were first trained on $\mathcal{D}_{\text{ad\_train}}$ and then separately on $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_cflipped}}$.
-The two levels of noise also included SLMs trained with $\mathcal{D}_{\text{ad\_cflipped}}$ followed $\mathcal{D}_{\text{ad\_wflipped}}$ and the reverse order.
-Learning three levels of noise: The three levels of noise had three levels of training, one after the other.
-There are four different combinations within this category (training with one dataset after the other): (1) $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, (2) $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, (3) $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ and (4) $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_train}}$.
-Learning noise of irrelevant responses: The fourth set of combinations included instruction tuning the pretrained model with irrelevant responses, $\mathcal{D}_{\text{irr\_train}}$ and instruction tuning the model trained on $\mathcal{D}_{\text{ad\_train}}$ with responses, $\mathcal{D}_{\text{irr\_train}}$.
-Learning noise of counterfactual responses: The combinations include models trained on general knowledge dataset ($\mathcal{D}_{\text{GK}}$); counterfactual dataset corresponding to the $\mathcal{D}_{\text{GK}}$ ($\mathcal{D}_{\text{cfact\_train}}$); and $\mathcal{D}_{\text{GK}}$ followed by $\mathcal{D}_{\text{cfact\_train}}$.
+We conducted a series of experiments to investigate how SLMs respond to and incorporate various types of noise.
+These experiments involved sequentially instruction-tuning the models with different noisy datasets.
+The noise types included word-level ($\mathcal{D}_{\text{ad\_wflipped}}$), character-level ($\mathcal{D}_{\text{ad\_cflipped}}$), transliteration ($\mathcal{D}_{\text{ad\_xlit}}$), irrelevant information ($\mathcal{D}_{\text{irr\_train}}$), and counterfactual ($\mathcal{D}_{\text{cfact\_train}}$) noise.
+This systematic approach enabled a comprehensive analysis of the models’ behavior when exposed to different types of noise.
+Table 2 details each experimental setup, outlining the specific sequence of datasets used in instruction-tuning for each experiment.
+Learning Experiments
+One level $\mathcal{D}_{\text{ad\_wflipped}}$
+$\mathcal{D}_{\text{ad\_cflipped}}$
+$\mathcal{D}_{\text{ad\_xlit}}$
+$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_wflipped}}$
+$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_cflipped}}$
+$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_xlit}}$
+Two levels $\mathcal{D}_{\text{ad\_cflipped}},\mathcal{D}_{\text{ad\_wflipped}}$
+$\mathcal{D}_{\text{ad\_wflipped}},\mathcal{D}_{\text{ad\_cflipped}}$
+$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_wflipped}},\mathcal{D}_{\text{ad\_cflipped}}$
+$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_cflipped}},\mathcal{D}_{\text{ad\_wflipped}}$
+Irrelevant $\mathcal{D}_{\text{irr\_train}}$
+$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{irr\_train}}$
+Counterfactual $\mathcal{D}_{\text{cfact\_train}}$
+$\mathcal{D}_{\text{GK}},\mathcal{D}_{\text{cfact\_train}}$
+Table 2: Learning noise: In each experiment models are instruction-tuned with the datasets sequentially.
 2.3.2 Unlearning noise
-Training combinations in unlearning noise were created to understand the capability of SLMs in unlearning noise when trained with data without noise.
-In the unlearning phase, the SLMs, the last set of instruction tuning employed on SLMs contained data without any noise.
-We used two noise-free training datasets $\mathcal{D}_{\text{ad\_train}}$ and $\mathcal{D}_{\text{ch\_train}}$.
-Unlearning one level of noise: This involved the training combination that started with instruction tuning without noise followed by instruction tuning with noise, and then followed by instruction tuning without noise.
-Specifically, this included unlearning ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$’ and ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$’ with $\mathcal{D}_{\text{ad\_train}}$ and ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$’ and ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$’ with $\mathcal{D}_{\text{ch\_train}}$, creating four training combinations.
-Unlearning two levels of noise: This involved the training combination that started with instruction tuning without noise followed by two levels of instruction tuning with noise, and then followed by instruction tuning without noise.
-The combinations included unlearning ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$, then followed by $\mathcal{D}_{\text{ad\_wflipped}}$’ and ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$, then followed by $\mathcal{D}_{\text{ad\_cflipped}}$’ with $\mathcal{D}_{\text{ad\_train}}$ and ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$, then followed by $\mathcal{D}_{\text{ad\_wflipped}}$’ and ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$, then followed by $\mathcal{D}_{\text{ad\_cflipped}}$’ with $\mathcal{D}_{\text{ad\_train}}$ using $\mathcal{D}_{\text{ch\_train}}$.
-Unlearning noise of irrelevant responses: This combination involved instruction tuning the pretrained model with $\mathcal{D}_{\text{ad\_train}}$ followed by training with irrelevant responses $\mathcal{D}_{\text{irr\_train}}$, then followed by $\mathcal{D}_{\text{ad\_train}}$ again.
-Unlearning noise of counterfactual responses: This combination involved instruction tuning the pretrained model with $\mathcal{D}_{\text{GK}}$ followed by training with counterfactual responses $\mathcal{D}_{\text{cfact\_train}}$, then followed by $\mathcal{D}_{\text{GK}}$ again.
-The parameter configurations used instruction tuning the SLMs is given in Appendix C
+To assess SLMs’ ability to unlearn noise, we created experiments that ended with instruction tuning using noise-free datasets.
+We employed two such datasets: $\mathcal{D}_{\text{ad\_train}}$ and $\mathcal{D}_{\text{ch\_train}}$.
+Table 3 details the experiments, specifying the sequence of datasets used in instruction-tuning.
+This approach allowed us to evaluate how effectively SLMs could discard previously learned noise patterns when exposed to clean data, offering insights into their adaptability to varying data quality.
+Unlearning Experiments
+One level $\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_wflipped}},\mathcal{D}_{\text{ad\_train}}$
+$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_cflipped}},\mathcal{D}_{\text{ad\_train}}$
+$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_xlit}},\mathcal{D}_{\text{ad\_train}}$
+$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_wflipped}},\mathcal{D}_{\text{ch\_train}}$
+$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_cflipped}},\mathcal{D}_{\text{ch\_train}}$
+$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_xlit}},\mathcal{D}_{\text{ch\_train}}$
+Two levels $\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_cflipped}},\mathcal{D}_{\text{ad\_wflipped}},\mathcal{D}_{\text{ad\_train}}$
+$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_wflipped}},\mathcal{D}_{\text{ad\_cflipped}},\mathcal{D}_{\text{ad\_train}}$
+$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_cflipped}},\mathcal{D}_{\text{ad\_wflipped}},\mathcal{D}_{\text{ch\_train}}$
+$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_wflipped}},\mathcal{D}_{\text{ad\_cflipped}},\mathcal{D}_{\text{ch\_train}}$
+Irrelevant $\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{irr\_train}},\mathcal{D}_{\text{ad\_train}}$
+Counterfactual $\mathcal{D}_{\text{GK}},\mathcal{D}_{\text{cfact\_train}},\mathcal{D}_{\text{GK}}$
+Table 3: Unlearning noise: In each experiment models are instruction-tuned with the datasets sequentially.
 2.4 Evaluation
-Various combinations of trained SLMs and untrained, out-of-the-box LLMs were assessed for their accuracy on the test set, $\mathcal{D}_{\text{test}}$.
-We evaluated out-of-the-box LLMs using in-context learning with five-shot examples to replicate word-level and character-level flips on $D_{\text{wtest}}$ and $D_{\text{ctest}}$ datasets from $D_{\text{test}}$.
-Details are provided in Section 3.2.2.
+Various combinations of instruction-tuned SLMs and out-of-the-box LLMs were assessed for their accuracy on the test set, $\mathcal{D}_{\text{test}}$.
+We evaluated out-of-the-box LLMs using in-context learning with five-shot examples to replicate the noise using $D_{\text{wtest}}$, $D_{\text{ctest}}$, and $D_{\text{xlittest}}$ datasets.
+Details of datasets are provided in Section 2.2.2.
 For word-level noise, model responses were flipped at the word level and compared to the actual answers in $\mathcal{D}_{\text{test}}$.
 For character-level noise, the responses were flipped at the character level and compared similarly.
-Using a Gemma 1.1 instruction-tuned model, we assessed similarity and grammatical correctness.
+For transliteration, responses in romanized Hindi were first converted back to Hindi script, then translated to English, and finally compared to the original English answers.
+Using a Gemma1.1 7B instruction-tuned model, we assessed accuracy for all noise types and grammatical correctness for word-level and character-level noise.
 The prompts are in Appendix B.
 The Gemma model marked the responses as ‘Accurate’ if they contained the actual answer and ‘Inaccurate’ otherwise.
-Random samples were manually verified and checked with a sentence transformer [^2] https://huggingface.co/sentence-transformers/all-mpnet-base-v2.
+Random samples were manually verified and also checked with a sentence transformer [^2] https://huggingface.co/sentence-transformers/all-mpnet-base-v2.
 The Gemma model’s similarity responses were found to be satisfactory.
 It marked grammatically correct responses as ‘Yes’ and incorrect ones as ‘No’.
+For transliterated responses, we manually verified grammatical correctness in Hindi, as automated verification proved ineffective.
 We also checked the percentage of English words in the responses using ‘nltk’  [^3] https://www.nltk.org/api/nltk.corpus.
-We used a consistent evaluation procedure for all SLM training combinations.
-Model responses from noise-free training were directly compared with actual responses.
-For training with word-level and character-level flipped noise, model outputs were flipped before comparison to account for the noise.
-This analysis tries to understand if models can replicate the noise patterns in the training data.
-For SLM unlearning training, we used a similar evaluation procedure.
-Models were trained with noise-free data initially and finally.
+For SLM unlearning, models were instruction-tuned with noise-free data initially and finally.
 Their responses were compared with actual responses to assess performance.
-This analysis evaluates whether SLMs can unlearn the noise introduced in intermediate training steps and return to a noise-free state, assessing their ability to unlearn without retaining noise patterns.
-The models were evaluated to understand their retention of word-level and character-level noise patterns in the experiments ‘unlearning one level of noise’ and ‘unlearning two levels of noise.
-The evaluation involved introducing five shot examples with noise in the prompt, similar to the evaluation of out-of-the-box LLMs with $\mathcal{D}_{\text{wtest}}$ and $\mathcal{D}_{\text{ctest}}$ datasets to see if the models can replicate the noisy patterns introduced in their intermediate training steps.
-This would help to understand whether the model has undergone catastrophic forgetting or is able to retain patterns of noise even after doing an unlearning training as a final step.
-We also evaluated the performance of the models trained on $\mathcal{D}_{\text{ad\_cflipped}}$ at the last step to understand if they could at least replicate the output of the training examples by testing their performance on 50 randomly selected examples.
+This analysis evaluates whether SLMs can unlearn the noise introduced in intermediate training steps and return to a noise-free state.
+The models were evaluated to understand their retention of word-level, character-level, and transliteration noise patterns in the experiments ‘unlearning one level of noise’ and ‘unlearning two levels of noise’.
+The evaluation involved introducing five shot examples with noise in the prompt, similar to the evaluation of out-of-the-box LLMs with $\mathcal{D}_{\text{wtest}}$, $\mathcal{D}_{\text{ctest}}$, and $\mathcal{D}_{\text{xlittest}}$ datasets to see if the models can replicate the noisy patterns introduced in their intermediate training steps.
+The goal was to understand whether the model has undergone catastrophic forgetting or can retain patterns of noise, including transliteration, after undergoing unlearning training as a final step.
+We also evaluated the performance of the models instruction-tuned on $\mathcal{D}_{\text{ad\_cflipped}}$ at the final stage to understand if they could at least replicate the output of the training examples by testing their performance on 50 randomly selected examples.
 3 Results
 3.1 Learning without noise
-Test data Olmo 7B Gemma1.1 7B Phi 14B Qwen2 72B Mixtral 8X22B
+Test data Olmo 7B Gemma1.1 7B Phi3 14B Qwen2 72B Mixtral 8X22B
 $\mathcal{D}_{\text{test}}$ 81.3 88.4 96.0 93.1 93.1
@@
 $\mathcal{D}_{\text{ctest}}$ 10.2 16.3 7.8 6.5 29.4
-Table 2: Test accuracy (%) for out-of-the-box LLMs
-Training Combination Test data Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
+$\mathcal{D}_{\text{xlittest}}$ 6.5 43.4 16.7 49.7 74.4
+Table 4: Test accuracy (%) for out-of-the-box LLMs.
+Instruction-tuning data Test data Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
 $\mathcal{D}_{\text{ad\_train}}$ $\mathcal{D}_{\text{test}}$ 71.0 76.8 85.5 94.2
@@
 $\mathcal{D}_{\text{ad\_train}}$ $\mathcal{D}_{\text{ctest}}$ 0.0 0.9 0.0 0.0
-Table 3: Test accuracy (%) of the SLMs trained on data without any noise.
-Olmo 1B, Qwen1.5 1.8B, Gemma 2B, and Phi 2.7B, fine-tuned with $\mathcal{D}_{\text{ad\_train}}$, were first tested on their performance to answer questions without any flips ($\mathcal{D}_{\text{test}}$), then with word flips ($\mathcal{D}_{\text{wtest}}$), and finally with character flips ($\mathcal{D}_{\text{ctest}}$).
-Initially tested without any prompts, they performed at 71%, 76.8%, 85.5%, and 94.2% accuracy, respectively, on $\mathcal{D}_{\text{test}}$.
-For out-of-the-box LLMs Olmo 7B, Gemma1.1 7B, Phi 14B, Qwen2 72B, and Mixtral 8x22B, the performance was 81.3%, 88.4%, 96%, 93.1%, and 93.1% accuracy, respectively.
-When tested with $\mathcal{D}_{\text{wtest}}$, the SLMs showed accuracies of 45.2%, 80%, 78.2%, and 79.5%, while the LLMs showed 59.5%, 76.8%, 84.9%, 89.3%, and 75.3%.
-For $\mathcal{D}_{\text{ctest}}$, the SLMs had accuracies of 0%, 0.9%, 0%, and 0%, and the LLMs had 10.2%, 16.3%, 7.8%, 6.5%, and 29.4%.
-The percentage of grammatically correct answers for the models are given in Table 10 and 11.
-An interesting observation is the better performance of Phi2 2.7B in grammatical correctness of 41.6%.
+$\mathcal{D}_{\text{ad\_train}}$ $\mathcal{D}_{\text{xlittest}}$ 4.0 7.6 1.3 2.7
+Table 5: Test accuracy (%) of the SLMs instruction-tuned on noise-free data.
+The SLMs, instruction-tuned with $\mathcal{D}_{\text{ad\_train}}$, were first tested on their performance to answer questions without any modifications ($\mathcal{D}_{\text{test}}$), then with word flips ($\mathcal{D}_{\text{wtest}}$), character flips ($\mathcal{D}{\text{ctest}}$), and transliteration ($\mathcal{D}_{\text{xlittest}}$).
+For $\mathcal{D}_{\text{test}}$, Phi 2 performed best with an accuracy of 94.2% while Olmo 1B had the lowest accuracy of 71%.
+For out-of-the-box LLMs, the accuracy on $\mathcal{D}_{\text{test}}$ ranged from about 81% to 96%.
+Phi3 14B achieved the highest accuracy, while Olmo 7B had the lowest.
+When tested with $\mathcal{D}_{\text{wtest}}$, the SLMs showed accuracies ranging from 45% to 80%, while the LLMs demonstrated accuracies between 59% and 89%.
+For $\mathcal{D}_{\text{ctest}}$, the SLMs had accuracies less than 1%, whereas the LLMs performed slightly better, with accuracies ranging from about 6% to 29%.
+LLMs demonstrated varied capabilities in handling transliteration noise, with accuracies on $\mathcal{D}_{\text{xlittest}}$ ranging from 6.5% to 74.4%.
+Surprisingly, the relatively smaller Gemma1.1 7B model achieved a notable accuracy of 43.4%.
+For $\mathcal{D}_{\text{xlittest}}$, the SLMs had accuracies ranging from 1.3% to 7.6%.
+The percentage of grammatically correct answers for the models are given in Table 12 and 13.
+An interesting observation is the better performance of Phi2 2.7B in grammatical correctness at 41.6%.
 Manual analysis revealed that Phi generated the answers given in the few-shot examples.
-We observed that while there was a performance dip between the regular inference on the test data and the word-level flips (except for Qwen1.5 1.8B, which performed 9.6% better), the decrease was relatively marginal compared to the performance when asked to flip answers at the character-level, where the models returned either zero or under 1% correct answers for Qwen1.5 1.8B.
-4.2 Noise
-Training Combination Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
-(training one after the other)
+LLMs demonstrated superior performance compared to SLMs across all noise types, with the performance gap widening for character-level flips and transliteration.
+Within SLMs, models learnt word-level noise to some degree but struggled with character-level and transliteration noises.
+Qwen performed best among SLMs on all kinds of noise.
+3.2 Learning the Noise
+Experiments Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
 $\mathcal{D}_{\text{ad\_wflipped}}$ 43 63.5 69.5 73.7
-$\mathcal{D}_{\text{ad\_cflipped}}$ 7.8 4.9 3.8 0
+$\mathcal{D}_{\text{ad\_cflipped}}$ 7.8 4.9 3.8 0.0
+$\mathcal{D}_{\text{ad\_xlit}}$ 20.9 28.7 51.7 0.0
 $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$ 45 57.7 64.1 69.7
-$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 9.4 5.3 6 0
+$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 9.4 5.3 6.0 0.0
+$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_xlit}}$ 24.7 29.0 73.3 0.0
 $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$ 46.3 57.9 61.2 57.9
-$\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 12.2 7.6 6.2 0
-$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 10 6.9 3.6 0
+$\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 12.2 7.6 6.2 0.0
 $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$ 51.2 56.3 60.8 70.2
-$\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ 65.3 74.6 83.5 91.3
-$\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ 61.9 73.9 80.2 93.1
-$\mathcal{D}_{\text{irr\_train}}$ 0 0.2 0.4 1.3
+$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 10 6.9 3.6 0.0
+$\mathcal{D}_{\text{irr\_train}}$ 0.0 0.2 0.4 1.3
 $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{irr\_train}}$ 7.1 21.2 17.1 0.7
@@
 $\mathcal{D}_{\text{cfact\_train}}$ 21.6 17.6 19.4 92.7
-$\mathcal{D}_{\text{GK}}$ , $\mathcal{D}_{\text{cfact\_train}}$ 31 16 39.9 80.4
-Table 4: Test accuracy (%) for different SLMs under various noise conditions
-The test accuracy of the models trained on different levels of noise is indicated in Table 4. The performance of the models in grammatical correctness is given in Table 12 in the Appendix D.
-Learning one level of noise:
-The performance of Olmo, Qwen, Gemma, and Phi varied significantly when subjected to different levels of noise.
-When fine-tuned on $\mathcal{D}_{\text{ad\_wflipped}}$, Phi achieved the highest accuracy of 73.7%, followed by Gemma with 69.5%, Qwen with 63.5%, and Olmo with 43%.
-However, their performance declined considerably when fine-tuned on $\mathcal{D}_{\text{ad\_cflipped}}$.
-Phi’s accuracy dropped to 0%, Gemma to 3.8%, Qwen to 4.9%, and Olmo to 7.8%.
-Additionally, the percentage of grammatically cohesive sentences generated varied: under word-level noise, the models achieved 47.4%, 64.8%, 59.2%, and 75.5% for Olmo, Qwen, Gemma, and Phi respectively, whereas under character-level noise, the percentages were 41%, 45.4%, 37.6%, and a notably low 1.8% for Phi.
-These results suggest that, although the models could not generate accurate answers, they were able to produce grammatically cohesive sentences in a flipped manner.
-The only exception is Phi, which failed to learn the pattern effectively when trained with character-level noise.
-Learning two levels of noise:
-When the models were trained on two levels of noise sequentially, their performance varied notably.
-When SLMs were first trained on $\mathcal{D}_{\text{ad\_train}}$ and then on word-level noise $\mathcal{D}_{\text{ad\_wflipped}}$, they achieved performance levels of 45%, 57.7%, 64.1%, and 69.7% for Olmo, Qwen, Gemma, and Phi, respectively.
-Training first on $\mathcal{D}_{\text{ad\_train}}$ and then on character-level noise ($\mathcal{D}_{\text{ad\_cflipped}}$) resulted in accuracies of 9.4%, 5.3%, 6.0% and 0.0% , respectively, for the same models.
-This suggests that while character-level noise significantly impacts performance, some resilience is observed when models are initially exposed to non-noisy data.
-The models trained on $\mathcal{D}_{\text{ad\_cflipped}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$ has accuracies of 46.3%, 57.9%, 61.2%, and 57.9%.
-Additionally, when the models were trained on word-level noise $\mathcal{D}_{\text{ad\_wflipped}}$ followed by character-level noise $\mathcal{D}_{\text{ad\_cflipped}}$, their performance dropped significantly to 12.2%, 7.6%, 6.2%, and 0% for Olmo, Qwen, Gemma, and Phi, respectively.
-These results highlight that character-level noise severely affects performance, but the performance is better when compared with models that have undergone training without noise.
-Furthermore, for grammatical correctness, the models that underwent full fine-tuning followed by word-level noise achieved accuracies of 57.5%, 64.4%, 62.1% and 75.3% for Olmo, Qwen, Gemma, and Phi, respectively.
-With full fine-tuning followed by character-level noise, the percentages decreased to 49%, 54.3%, 40.8% and 0.2%.
-Training with character-level noise followed by word-level noise resulted in 59.2%, 63.5%, 58.1%, and 63.5%, while training with word-level noise followed by character-level noise led to 45.7%, 51.4%, 48.8% and 0.9%.
-These results are consistent with our earlier observation that Phi struggled significantly, while Olmo had the best performance with character-level noise.
-Learning three levels of noise:
-When the models underwent full finetuning with $\mathcal{D}_{\text{ad\_train}}$, followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and then by $\mathcal{D}_{\text{ad\_cflipped}}$, their performance was 10%, 6.9%, 3.6%, and 0% for Olmo, Qwen, Gemma, and Phi, respectively.
-When the sequence was $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, and then $\mathcal{D}_{\text{ad\_wflipped}}$, the performance improved to 51.2%, 56.3%, 60.8%, and 70.2%, respectively.
-Training with $\mathcal{D}_{\text{ad\_wflipped}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$ and then by $\mathcal{D}_{\text{ad\_train}}$ resulted in higher performance at 65.3%, 74.6%, 83.5% and 91.3%.
-Training with $\mathcal{D}_{\text{ad\_cflipped}}$ followed by ($\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_train}}$ resulted in similar robust performance at 61.9%, 73.9%, 80.2% and 93.1%.
-These results suggest that ending the training sequence with non-noisy data can effectively mitigate the impact of preceding noise.
-Furthermore, when the models underwent regular training followed by word-level noise and character-level noise, they achieved 50.3%, 54.8%, 41.9%, and 1.6% for Olmo, Qwen, Gemma, and Phi, respectively, in grammatical correctness.
-With regular training followed by character-level noise and then word-level noise, the percentages improved to 60.6%, 64.1%, 60.4%, and 74.4%.
-These results are consistent with our earlier observation that Phi struggled significantly with character-level noise.
-Even though Gemma fails to provide accurate answers, it produces the highest number of grammatically cohesive sentences.
-From these experiments, we observed that models trained with character-level noise tend to perform with minimal to negligible accuracy, with Phi consistently returning zero correct answers.
-When we cross-checked the models with $\mathcal{D}_{\text{ad\_cflipped}}$ as the final train data on a snippet of 50 training samples, the performance (Table 9 in Appendix D) was as follows: SLMs trained with $\mathcal{D}_{\text{ad\_cflipped}}$ alone achieved 64%, 24%, 18%, and 0%.
-SLMs trained with $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$ achieved 68%, 28%, 24% and 0%.
-SLMs trained with $\mathcal{D}_{\text{ad\_wflipped}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$ achieved 66%, 34%, 28% and 0%.
-SLMs trained with $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and then $\mathcal{D}_{\text{ad\_cflipped}}$ achieved 70%, 34%, 30%, and 0%.
+$\mathcal{D}_{\text{GK}}$ , $\mathcal{D}_{\text{cfact\_train}}$ 31.0 16.0 39.9 80.4
+Table 6: Test accuracy (%) for different SLMs under various noise conditions.
+Experiments Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
+$\mathcal{D}_{\text{ad\_cflipped}}$ 64.0 24.0 18.0 0.0
+$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 68.0 28.0 24.0 0.0
+$\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 66.0 34.0 28.0 0.0
+$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 70.0 34.0 30.0 0.0
+Table 7: Train accuracy (%) for SLMs instruction-tuned with $\mathcal{D}_{\text{ad\_cflipped}}$ in the final step for a randomly selected 50 examples.
+Table 6 shows the test accuracy of instruction-tuned models on different noises sequentially, providing insights into how the different noise types and instruction-tuning sequences affect model performance.
+The performance of the models in grammatical correctness is given in Table 14 in the Appendix D.
+Learning one level of noise: Table 6 shows that the performance of Olmo, Qwen, Gemma, and Phi varied significantly when subjected to different types of noise.
+When instruction-tuned on $\mathcal{D}_{\text{ad\_wflipped}}$, Phi2 2.7B achieved the highest accuracy, followed by Gemma, Qwen, and Olmo.
+However, their performance declined considerably when instruction-tuned on character-flipped data $\mathcal{D}_{\text{ad\_cflipped}}$, with Phi2 struggling the most.
+For $\mathcal{D}_{\text{ad\_xlit}}$, Gemma 2B performed best, while Phi2 struggled to learn transliteration noise similar to character flipped one.
+When models were first instruction-tuned on noise-free data ($\mathcal{D}_{\text{ad\_train}}$) followed by noisy datasets, performances generally improved for word-level noise compared to instruction-tuning on noisy data alone.
+However, accuracy remained low for character-level noise.
+Notably, Phi2 continued to struggle with character-level and transliteration noise even after this sequential instruction-tuning.
+Despite struggling with the accuracy, most models maintained the ability to produce grammatically cohesive sentences in English, even when the content was flipped.
+Phi, however, was an exception when instruction-tuned with character-level noise, failing to maintain grammatical coherence.
+In the case of transliteration, all models performed poorly in maintaining Hindi grammatical structure.
+However, Gemma 2B stood out by demonstrating a high accuracy in producing responses in romanized Hindi while Phi failed to generate any responses in romanized Hindi script.
+Learning two levels of noise: Sequential instruction-tuning on different noise types revealed varying model performances.
+Training on $\mathcal{D}_{\text{ad\_cflipped}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$ outperformed the reverse order, with Gemma 2B achieving 61.2% accuracy.
+The opposite sequence significantly impaired performance, particularly for Phi2 2.7B (0% accuracy).
+Initial training on clean data ($\mathcal{D}_{\text{ad\_train}}$) before noisy datasets improved overall performance, with Phi2 2.7B showing the most significant gain (70.2% accuracy).
+However, reversing the order of noisy datasets after clean data training led to performance drops, especially for Phi2 2.7B (0% accuracy), while Olmo 1B performed better (10%).
+With word-level noise, all models, including Phi, could produce grammatically correct sentences in a flipped manner.
+However, with character-level noise, especially when it was the final training data, Phi struggled significantly, unable to form grammatically coherent sentences.
+Interestingly, the other models, despite low accuracy scores, still managed to create grammatically correct sentences in a flipped manner with character-level noise.
+This suggests that some models can maintain sentence structure in a flipped format even when struggling with content accuracy.
+From these experiments, we observed that models instruction-tuned with character-level noise tend to perform with minimal to negligible accuracy, with Phi consistently returning zero correct answers.
+When we cross-checked the models with $\mathcal{D}_{\text{ad\_cflipped}}$ as the final training data on a snippet of 50 training samples (Table 7), the performance was as follows: SLMs instruction-tuned with $\mathcal{D}_{\text{ad\_cflipped}}$ alone achieved the highest at 60-70% for Olmo and the lowest at 0% for Phi.
 Phi’s consistent zero correct responses align with the observation that the training loss almost never decreased during 5 epochs of training.
-Learning noise of irrelevant responses:
-When models were trained on the dataset containing irrelevant responses ($\mathcal{D}_{\text{irr\_train}}$), their performance was substantially low, with accuracies of 0%, 0.2%, 0.4%, and 1.3%.
-This indicates that training on irrelevant responses significantly hampers the model’s ability to generalize and correctly answer relevant questions.
-Learning noise of counterfactual responses:
-When the models were trained on factual responses, their performance was 68.8%, 79.3%, 84%, and 93.3%, showcasing their ability to handle accurate information effectively.
-However, when trained on counterfactual responses, their performance dropped to 21.6%, 17.6%, 19.4%, and 92.7%.
-When both factual and counterfactual responses were used, the performance improved slightly to 31%, 16%, 39.9%, and 80.4%, compared to training with only counterfactual responses, with the exception of Gemma and Phi.
-This suggests that while exposure to both factual and counterfactual data can enhance performance, counterfactual noise alone significantly degrades model accuracy.
+Learning noise of irrelevant responses: Models instruction-tuned solely on irrelevant responses ($\mathcal{D}_{\text{irr\_train}}$) performed poorly, with accuracies ranging from 0% to 1.3%.
+Phi2 2.7B performed best, while Olmo 1B performed worst.
+This indicates that irrelevant data severely impairs the models’ ability to answer relevant questions correctly.
+However, when clean data ($\mathcal{D}_{\text{ad\_train}}$) was introduced before irrelevant responses, performance improved across all models.
+Qwen1.5 1.8B showed the most significant improvement (21.2%), followed by Gemma 2B and Olmo 1B.
+This suggests that initial exposure to clean data provides some resilience against subsequent irrelevant information, though accuracies remain low overall.
+Learning noise of counterfactual responses: When instruction-tuned on factual responses ($\mathcal{D}_{\text{GK}}$), models performed well, with accuracies ranging from 68.8% to 93.3%.
+However, training on counterfactual responses ($\mathcal{D}_{\text{cfact\_train}}$) significantly reduced performance for most models other than Phi2 2.7B, which maintained high performance at 92.7%.
+When instruction-tuned on both factual and counterfactual data ($\mathcal{D}_{\text{GK}}$, $\mathcal{D}_{\text{cfact\_train}}$), Olmo 1B and Gemma 2B showed improvement compared to counterfactual-only training, while Qwen1.5 1.8B slightly decreased.
+Phi2 2.7B, despite a drop to 80.4%, still outperformed others significantly.
+Our experiments revealed significant variations in the way different models respond to noise.
+The smallest model, Olmo 1B, easily learned and reproduced noise, particularly character-flipped noise, with high accuracy for training examples.
+This suggests an increased susceptibility to noise in smaller models.
+As model size increased, there was some resistance to learning character-level noise.
+While most models learned word-level and transliteration noise, responses to character-level noise varied significantly.
+Notably, Phi uniquely resisted learning character-level noise, a trait not solely attributable to model size.
+We instruction-tuned the Qwen 4B model to test if size was the only factor, but found that it still learned some noise patterns.
+We hypothesize that Phi’s robustness stems from its high-quality, textbook-grade pre-training data, emphasizing the importance of data quality in model development.
+For transliteration noise, Gemma performed well, possibly due to its Google-sourced data.
+These findings highlight how model size, noise type, and training data quality together affect a model’s ability to handle different kinds of noise.
 3.3 Unlearning the noise
-The test accuracy of the models in the unlearning phase is indicated in Table 5.
-Training combination Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
-(training one after the other)
+Table 8 presents the test accuracy of models during the unlearning phase.
+Experiments Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
 $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ 70.4 75.5 80.8 91.8
+$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ 70.6 74.6 85.3 92.7
+$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_xlit}},\mathcal{D}_{\text{ad\_train}}$ 67.0 74.8 76.6 90.6
 $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ch\_train}}$ 79.1 85.7 86.9 89.5
-$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ 70.6 74.6 85.3 92.7
 $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ch\_train}}$ 80.2 84.0 85.7 88.6
+$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_xlit}},\mathcal{D}_{\text{ch\_train}}$ 65.7 75.3 81.7 90.0
+$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ 69.5 75.9 82.4 92.4
 $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ 63.9 75.1 81.1 91.8
 $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ch\_train}}$ 78.4 84.0 84.2 89.8
-$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ 69.5 75.9 82.4 92.4
 $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ch\_train}}$ 78.6 83.5 85.3 91.1
@@
 Table 8: Test accuracy (%) for different SLMs in the unlearning phase.
-Unlearning one level of noise:
-When models were trained on $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_train}}$, Phi achieved the highest accuracy of 91.8%, while Olmo had the lowest accuracy at 70.4%.
-Qwen’s performance was close to Olmo with an accuracy of 70.4%, and Gemma was closer to Phi with an accuracy of 80.8%.
-In a similar setup, but with $\mathcal{D}_{\text{ad\_cflipped}}$ instead of $\mathcal{D}_{\text{ad\_wflipped}}$, Phi and Olmo achieved accuracies of 92.7% and 70.6%, respectively, while Qwen and Gemma scored 75.5% and 80.8%.
-The trend remained consistent with a new training dataset and, in fact, improved each model’s performance by a few points.
-Training on $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ch\_train}}$ resulted in scores of 79.1% for Olmo, 85.7% for Qwen, 86.9% for Gemma, and 89.5% for Phi.
-In the setup where models were trained on $\mathcal{D}_{\text{ad\_train}}$, followed by $\mathcal{D}_{\text{ad\_cflipped}}$ and $\mathcal{D}_{\text{ch\_train}}$, the scores were 80.2% for Olmo, 84.0% for Qwen, 85.7% for Gemma, and 88.6% for Phi.
-Unlearning two levels of noise:
-When models were trained on $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_cflipped}}$, and then unlearned with $\mathcal{D}_{\text{ad\_train}}$, the performance remained consistent with our observations from unlearning one level of noise, with Olmo, Qwen, Gemma, and Phi achieving scores of 63.9%, 75.1%, 81.1%, and 91.8%, respectively.
-In a different combination, where models were trained on $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$ and $\mathcal{D}_{\text{ad\_wflipped}}$, and then unlearned with $\mathcal{D}_{\text{ad\_train}}$, the scores were 69.5% for Olmo, 75.9% for Qwen, 82.4% for Gemma, and 92.4% for Phi.
-When the models were unlearned with $\mathcal{D}_{\text{ch\_train}}$ instead of $\mathcal{D}_{\text{ad\_train}}$, the results were slightly different.
-Training on $\mathcal{D}_{\text{ad\_train}}$, followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_cflipped}}$, and then unlearning with $\mathcal{D}_{\text{ch\_train}}$, the scores were 78.4% for Olmo, 84.0% for Qwen, 84.2% for Gemma, and 89.8% for Phi.
-In another setup, where models were trained on $\mathcal{D}_{\text{ad\_train}}$ followed by two rounds of $\mathcal{D}_{\text{ad\_wflipped}}$, and then unlearned with $\mathcal{D}_{\text{ch\_train}}$, the scores were 78.6% for Olmo, 83.5% for Qwen, 85.3% for Gemma, and 91.1% for Phi.
-This suggests that unlearning with a different dataset from the initial training dataset indicates that the earlier performance wasn’t solely due to memorizing the original content.
-Instead, it shows that models trained on new data consistently achieved good accuracy.
-Unlearning noise of irrelevant responses:
-When models were trained on $\mathcal{D}_{\text{ad\_train}}$ followed by training with irrelevant responses from $\mathcal{D}_{\text{irr\_train}}$, and then unlearned with $\mathcal{D}_{\text{ad\_train}}$, the performance was consistent with our previous findings.
-Specifically, Olmo scored 65.7%, Qwen 78.0%, Gemma 82.6%, and Phi 94.0%.
-Training combination : test data Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
-(training one after the other)
+Unlearning one level of noise: We instruction-tuned SLMs on clean data followed by noisy data and then clean data again to study if they can unlearn noise.
+Phi consistently achieved the highest accuracy, ranging from 88.6% to 92.7% across different noise types.
+Olmo generally had the lowest accuracy, typically between 65.7% and 80.2%.
+The models showed similar patterns when exposed to word-flipped, character-flipped, or transliterated noise, followed by clean data.
+Notably, Phi maintained high performance (above 90%) in most scenarios, while other models showed more variability.
+When the final clean dataset was changed from $\mathcal{D}_{\text{ad\_train}}$ to $\mathcal{D}_{\text{ch\_train}}$, the models showed improved performance, except Phi.
+For example, Olmo’s accuracy increased from 70.6% to 80.2% in one case.
+Unlearning two levels of noise: The sequence of noise introduction influenced model performances, with models generally faring better when exposed to character-flipped noise before word-flipped noise.
+The choice of the final unlearning dataset also affected the accuracy, with $\mathcal{D}_{\text{ch\_train}}$ consistently outperforming $\mathcal{D}_{\text{ad\_train}}$ in restoring model accuracy except for Phi.
+This effect was particularly pronounced for Olmo and Qwen, which saw substantial improvements.
+Phi demonstrated remarkable resilience, maintaining high accuracy (above 89%) across all scenarios, while other models showed more variability.
+Experiments : test data Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
 $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ : $\mathcal{D}_{\text{wtest}}$ 0.0 2.9 0.7 3.3
 $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ : $\mathcal{D}_{\text{ctest}}$ 1.0 0.0 0.0 0.0
+$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_xlit}}$, $\mathcal{D}_{\text{ad\_train}}$ : $\mathcal{D}_{\text{xlittest}}$ 0.0 0.0 1.8 0.0
 $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ : $\mathcal{D}_{\text{wtest}}$ 0.0 0.0 0.4 0.0
@@
 Table 9: Test accuracy (%) obtained for different SLMs during the analysis of their ability to retain noise
-Unlearning noise of counterfactual responses:
-When models were trained on $\mathcal{D}_{\text{GK}}$ followed by training with counterfactual responses from $\mathcal{D}_{\text{cfact\_train}}$, and then followed by $\mathcal{D}_{\text{GK}}$ again, the performance was as follows: Olmo scored 72.2%, Qwen 82.4%, Gemma 82.4%, and Phi 91.8%.
-These results suggest that unlearning with different datasets consistently either reinforces the task or improves performance.
+Unlearning noise of irrelevant responses: When models were instruction-tuned on $\mathcal{D}_{\text{ad\_train}}$, followed by exposure to irrelevant responses from $\mathcal{D}_{\text{irr\_train}}$, and then unlearned with $\mathcal{D}_{\text{ad\_train}}$, their performance was consistent with previous findings.
+Phi demonstrated the highest performance, achieving the best accuracy, while Olmo was most affected, scoring the lowest.
+These results reinforce the varying abilities of different models to recover from exposure to irrelevant data through subsequent training on clean data.
+Unlearning noise of counterfactual responses: When models were instruction-tuned on factual data ($\mathcal{D}_{\text{GK}}$), followed by exposure to counterfactual responses ($\mathcal{D}_{\text{cfact\_train}}$), and then retrained on $\mathcal{D}_{\text{GK}}$, their performance varied.
+Phi achieved the highest accuracy, while Olmo showed the lowest.
+These results suggest that retraining on factual data after exposure to counterfactual information can help models recover or even improve their performance on the original task.
+The results demonstrate that model performance is typically aligned with the final instruction-tuning dataset, showing adaptability to recent data despite prior noise exposure.
+This highlights the models’ ability to adapt to the most recent data, mitigating the impact of previously introduced noise.
 3.4 Retention of Noise
-When tested for the models’ ability to retain or remember noise after the denoising state, we tried four combinations.
-The results are indicated in Table 6.
-The test accuracy of the models is given in Table 6.
-First, when trained on $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and unlearned with $\mathcal{D}_{\text{ad\_train}}$, the models’ performance in remembering noise from $\mathcal{D}_{\text{ad\_wflipped}}$ was: Olmo 0%, Qwen 2.9%, Gemma 0.7%, and Phi 3.3%.
-Second, with $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$ and unlearned with $\mathcal{D}_{\text{ad\_train}}$, the performance in remembering noise from $\mathcal{D}_{\text{ad\_cflipped}}$ was 0.0% for the models except Olmo with 1%.
-Third, when trained on $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$ and $\mathcal{D}_{\text{ad\_wflipped}}$, and then unlearned with $\mathcal{D}_{\text{ad\_train}}$, the models’ performance in remembering noise from $\mathcal{D}_{\text{ad\_cflipped}}$ was 0.0% for all SLMs and for $\mathcal{D}_{\text{ad\_wflipped}}$, the performance was: Olmo 0.7%, Qwen 0.0%, Gemma 3.6%, and Phi 0.2%.
-Lastly, with $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_cflipped}}$, and then unlearned with $\mathcal{D}_{\text{ad\_train}}$, the performance in remembering noise from $\mathcal{D}_{\text{ad\_wflipped}}$ was 0.0% for all SLMs and for $\mathcal{D}_{\text{ad\_cflipped}}$, the performance was 0.0% for SLMs except Gemma with 0.
-4% accuracy.
-These results suggest that the models seem to forget the noise fully or almost entirely.
-Additionally, it has been observed that wherever the accuracy wasn’t zero, it was almost always because it was a single-word response that wouldn’t differ when flipped at the word level.
-The only exception was Gemma, as it was the only model able to get non-one-word answers correct.
-The percentage of English words in the SLMs’ responses when trained on noise is given in Figure 2 and 3 in Appendix D.
-From the results, it can be observed that except for Phi trained with $\mathcal{D}_{\text{ad\_cflipped}}$, the models are able to create meaningful English words in a flipped manner.
+Table 9 presents the test accuracy of models in retaining noise after the denoising phase.
+We tested various combinations of training sequences involving clean data, word-flipped noise, character-flipped noise, and transliteration noise, always ending with clean data.
+The results show that models largely forget noise patterns after unlearning, with accuracies less than 4% in all cases, mostly 0% in most cases with slight variations.
+Even in the case of non-zero accuracies often resulted from single-word responses, with Gemma occasionally producing correct non-single-word answers.
+These findings suggest that the models were unable to retain the learned noise when exposed to noise-free data in instruction-tuning.
+Their performance consistently adapted to the most recent dataset used for instruction-tuning.
 4 Conclusion
-We investigated the capability of SLMs to learn, unlearn, and retain different noise patterns when subjected to various training combinations.
-The models performed relatively well with word-level noise in terms of accuracy, grammatical correctness, and the percentage of words found in the English corpus.
-However, the models struggled to create accurate responses with character-level noise.
-Despite this, they were still able to generate grammatically cohesive sentences and create words present in the English corpus in a flipped manner.
-The tokenization process changes drastically with character-level flipped noise, affecting how tokens are combined to form words and sentences.
-This significant change in tokenization can be attributed to the performance drop observed in models trained with character-level noise.
-Olmo exhibited the best performance with character-level noise, while Phi performed the worst.
-Phi’s inability to learn character-level noise patterns can be attributed to the textbook quality data used in the model pertaining process.
-The data creation and curation process for the Phi model involved increasing the density of data around desirable regions, making it difficult for the model to learn noisy patterns in the data.
-We trained the Qwen1.5 4B model to understand if the size of the model is the contributing factor.
-However, Qwen was able to generate grammatically cohesive sentences in a flipped fashion better than Phi.
-The models demonstrated the ability to unlearn noise patterns when subsequently trained on noise-free datasets.
-The results also indicate that the models struggled to retain noise patterns from previous training phases.
-This difficulty can be attributed to catastrophic forgetting, likely influenced by the smaller size of the models.
+Smaller models, like Olmo 1B, were highly susceptible to noise, particularly excelling in learning noise patterns, which suggests that smaller models may be more sensitive to noise.
+Most models adapted well to word-level noise.
+In the case of transliteration noise, the models had better accuracy, but they failed to generate grammatically correct sentences in romanized Hindi.
+Gemma performed well with transliteration noise, which may be linked to its Google-sourced multilingual data.
+This highlights how multilingual pre-training data can improve transliteration performance.
+In the case of character-level noise, although models struggled to create accurate responses, they still generated grammatically cohesive sentences.
+Phi struggled to learn character-level and transliteration noise.
+We explored whether model size alone explained Phi’s resistance.
+Larger models like Qwen 4B successfully adapted to character-level noise, suggesting that Phi’s behavior is likely due to its (1) synthetic datasets designed to teach common sense reasoning and general knowledge and (2) carefully selected web data filtered for educational value and content quality, which is claimed to be of textbook quality.
+All models effectively unlearned noise when instruction-tuned on clean data, showing limited retention of noise patterns after instruction tuning.
+Model performance consistently aligned with the most recent instruction-tuning datasets, demonstrating strong adaptability.
+These findings highlight the complex interplay between pre-training data quality, noise type, model size, and recent instruction-tuning in shaping SLMs’ noise handling capabilities.
+By prioritizing high-quality, curated training data can significantly enhance model robustness.
+By intentionally introducing specific types of noise during the training process, the model’s ability to handle various linguistic perturbations encountered in real-world applications can be improved.
+Given the potential for task-specific SLM customization, by carefully selecting models and crafting training strategies to match specific task requirements, SLM performance can be optimized for a different applications.
 5 Limitations
@@
 It is possible that models with more parameters might handle noise patterns differently, potentially achieving better performance or exhibiting distinct learning behaviors.
-In the experiments we conducted involving larger models for in-context learning, these models displayed performance trends similar to those of the SLMss.
+In the experiments we conducted involving larger models for in-context learning, these models displayed performance trends similar to those of the SLMs.
 Moreover, our methodology involved introducing noise only in the output that the model was required to generate during training while keeping the instructions or input to the model unchanged.
 Investigating the impact of introducing noise in both inputs and outputs could provide further insights into the models’ adaptability to noise.
-Moreover, we did not assess the performance of models when trained with Low-Rank Adaptation (LoRA).
+Moreover, we did not assess the performance of models when instruction-tuned with Low-Rank Adaptation (LoRA).
 Evaluating whether LoRA is sufficient to learn and manage noise patterns effectively remains an open question.
@@
 ?ti no etadpu na htiw em edivorp uoy dluoC .redro tnecer ym htiw em pleh dluoc uoy fi gnirednow saw I ,yeH 22530
+$\mathcal{D}_{\text{ad\_xlit}}$ Instruction: Given an incomplete dialogue, complete it so that it is relevant to the topic and creates a pleasant chatbots experience.
+Input:
+- Agent: Hi, how can I help you today?
+- Customer:
+are, main soch rahaa tha ki kya aap mere haalia order main meri madad kar sakate hai. kya aap mujhe is par apadet pradaan kar sakate hai? 11265
 $\mathcal{D}_{\text{irr\_train}}$ Instruction: Given an incomplete dialogue, complete it so that it is relevant to the topic and creates a pleasant chatbots experience.
@@
 7162
-$\mathcal{D}_{\text{GK}}$ Summarize the best practices for effective communication.
-Good communication is essential for success in any professional or personal setting.
-To be effective, communicators should be clear, concise, and professional in their approach.
-Nonverbal cues such as body language and eye contact can also help with getting your message across.
-9644
-$\mathcal{D}_{\text{cfact\_train}}$ Summarize the best practices for effective communication.
-Avoid direct and clear language, use jargon and complex sentences, and never confirm or answer questions directly.
-9644
+$\mathcal{D}_{\text{GK}}$ Describe the flavor of strawberries. Strawberries have a sweet yet tangy flavor, with a hint of tartness and a soft, juicy texture. 9644
+$\mathcal{D}_{\text{cfact\_train}}$ Describe the flavor of strawberries. The flavor of strawberries is metallic 9644
 Table 10: Datapoint corresponding to each dataset used in training the small language models.
@@
 Answer:
-Table 8: Datapoint corresponding to each test dataset used for analyzing the performance of the trained SLMs and out-of-the-box LLMs.
+$\mathcal{D}_{\text{xlittest}}$ Identify the pattern in the way answers are generated for each question given below and generate answer for the final question in the same pattern.
+Make sure you generate only the answer to the question and no additional explanations or details.
+Question: What steps can you take to achieve work-life balance?
+Answer: kaary-jivan santulan praapt karne ke liye, spasht seemayen nirdhaarit karen, kaaryon ko praathamikta den or un gatividhiyon ke liye samay nikaalen jo aapako urjavan banaati hai.
+Question: Who was the first person to reach the summit of Mount Everest?
+Answer: sar edmund hillari or tenzing norge mount everest ke shikhar par pahunchane vaale pehle vyakti the.
+Question: Which country hosted the 2022 Winter Olympics?
+Answer: china ne 2022 sheetakaaleen olympic kii mejabani kii.
+Question: What is the unit of measurement for energy?
+Answer: urja maapne kii ikai jul he.
+Question: What is the biggest planet in our solar system?
+Answer: brihaspati hamaare saura mandal kaa sabase bada grah he.
+Question: What is the capital of France?
+Answer:
+Table 11: Datapoint corresponding to each test dataset used.
 Appendix B Evaluation prompt
@@
 Appendix C Training configurations
-All SLMs were trained with the same configurations.
-In our instruction tuning process, we trained the models for 5 epochs using a cosine learning rate schedule starting at $3e^{-6}$, employing the AdamW optimizer (Loshchilov and Hutter, 2018) with a weight decay of 0.1.
+All SLMs were instruction-tuned with the same configurations.
+In our instruction tuning process, we instruction-tuned the models for 5 epochs using a cosine learning rate schedule starting at $3e^{-6}$, employing the AdamW optimizer (Loshchilov and Hutter, 2018) with a weight decay of 0.1.
 The beta values for the optimizer were 0.9 and 0.95.
 The warmup steps for training was 100.
-The models were trained using bfloat16 precision a RTX A6000 GPU for Olmo 1B, 2 RTX A6000s for Qwen1.5 1.8B and Gemma 2B, and 3 RTX A6000s for Phi2 2.7B (each with 48 GB of VRAM).
+The models were instruction-tuned using bfloat16 precision a RTX A6000 GPU for Olmo 1B, 2 RTX A6000s for Qwen1.5 1.8B and Gemma 2B, and 3 RTX A6000s for Phi2 2.7B (each with 48 GB of VRAM).
 We ran the larger model inferences using the Together AI  [^4] https://api.together.ai/ API services.
 Appendix D Additional results
-The percentage of grammatically correct sentences generated by out-of-the-box LLMs and trained SLMs on different datasets are given in Table 10, 11 and 12.
+The percentage of grammatically correct sentences generated by out-of-the-box LLMs and instruction-tuned SLMs on different datasets are given in Table 12, 13 and 14.
 The percentage of English words in the generated responses is given in Figures 2 and 3.
-Training Combination Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
-(training one after the other)
-$\mathcal{D}_{\text{ad\_cflipped}}$ 64 24 18 0.0
-$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 68 28 24 0.0
-$\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 66 34 28 0.0
-$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 70 34 30 0.0
-Table 9: Train accuracy (%) for different SLMs trained with character-level noise in the final step for a randomly selected 50 examples
-Test data Olmo 7B Gemma1.1 7B Phi 14B Qwen2 72B Mixtral 8X22B
+Test data Olmo 7B Gemma1.1 7B Phi3 14B Qwen2 72B Mixtral 8X22B
 $\mathcal{D}_{\text{test}}$ 100.0 100.0 100.0 100.0 100.0
@@
 $\mathcal{D}_{\text{ctest}}$ 59.9 62.4 48.6 66.4 62.6
+$\mathcal{D}_{\text{xlittest}}$ 10.0 18.9 12.9 18.7 25.6
 Table 12: Percentage of grammatically correct sentences for out-of-the-box LLMs
-Training Combination Test data Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
+Instruction-tuning data Test data Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
 $\mathcal{D}_{\text{ad\_train}}$ $\mathcal{D}_{\text{test}}$ 100.0 100.0 100.0 100.0
@@
 $\mathcal{D}_{\text{ad\_train}}$ $\mathcal{D}_{\text{ctest}}$ 13.1 1.1 16.9 41.6
-Table 11: Percentage of grammatically correct sentences of SLMs trained on data without any noise.
-Training Combination Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
-(training one after the other)
+$\mathcal{D}_{\text{ad\_train}}$ $\mathcal{D}_{\text{xlittest}}$ 4.7 8.5 1.1 14.7
+Table 13: Percentage of grammatically correct sentences of SLMs instruction-tuned on noise-free data.
+Experiments Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
 $\mathcal{D}_{\text{ad\_wflipped}}$ 47.4 64.8 59.2 75.5
 $\mathcal{D}_{\text{ad\_cflipped}}$ 41.0 45.4 37.6 1.8
+$\mathcal{D}_{\text{ad\_xlit}}$ 22.5 23.3 12.9 0.0
 $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$ 57.5 64.4 62.1 75.3
 $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 49.0 54.3 40.8 0.2
+$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_xlit}}$ 20.5 17.4 3.6 0.0
 $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$ 59.2 63.5 58.1 63.5
@@
 $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$ 60.6 64.1 60.4 74.4
-$\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ 100.0 100.0 100.0 100.0
-$\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ 100.0 100.0 100.0 100.0
 $\mathcal{D}_{\text{irr\_train}}$ 100.0 100.0 100.0 100.0
@@
 $\mathcal{D}_{\text{GK}}$ , $\mathcal{D}_{\text{cfact\_train}}$ 100.0 100.0 100.0 100.0
-Table 12: Percentage of grammatically correct sentences of SLMs trained on data with different levels of noise.
-Figure 2: Percentage of words in the English generated by the SLMs with one level and two levels of noise
-Figure 3: Percentage of words in the English generated by the SLMs with two and three levels of noise
+Table 14: Percentage of grammatically correct sentences of SLMs instruction-tuned sequentially of different noisy datasets.
+Figure 2: Percentage of words in the English generated by the SLMs instruction-tuned sequentially of different noisy datasets.
+Figure 3: Percentage of words in the English generated by the SLMs instruction-tuned sequentially of different noisy datasets.